### v2

# ConTextTab: A Semantics-Aware Tabular In-Context Learner

Abstract
Tabular in-context learning (ICL) has recently achieved state-of-the-art (SOTA) performance on several tabular prediction tasks. Previously restricted to classification problems on small tables, recent advances such as TabPFN [17] and TabICL [28] have extended its use to larger datasets. Although current table-native ICL architectures are architecturally efficient and well-adapted to tabular data structures, their exclusive training on synthetic data limits their ability to fully leverage the rich semantics and world knowledge contained in real-world tabular data. At the other end of the spectrum, tabular ICL models based on pretrained large language models such as TabuLa-8B [11] integrate deep semantic understanding and world knowledge but are only able to make use of a small amount of context due to inherent architectural limitations. With the aim to combine the best of both these worlds, we introduce ConTextTab, integrating semantic understanding and alignment into a table-native ICL framework. By employing specialized embeddings for different data modalities and by training on large-scale real-world tabular data, our model is competitive with SOTA across a broad set of benchmarks while setting a new standard on the semantically rich CARTE benchmark. Code and checkpoints are available at: https://github.com/SAP-samples/contexttab.
1 Introduction
Tables with information spread across rows and columns remain a predominant data format in many real-world applications [3], making their understanding through machine learning algorithms critical. Despite the great success of deep learning approaches in natural language processing and computer vision, leveraging large amounts of pretraining data, conventional machine learning methods such as gradient boosted trees and their variants remain the predominant state-of-the-art (SOTA) across tabular prediction benchmarks. Recently, however, applying the in-context learning (ICL) paradigm to tabular prediction tasks has shown promising results by enabling pretraining of deep learning models across a large set of heterogeneous tables, constituting a new SOTA on small to medium tabular prediction tasks [17]. In this setting, predicting a target value $y$ based on the input features $\boldsymbol{x}$ of a row in a table $T$ additionally uses further rows from $T$ (the context), including their target values. This enables the model to adapt to new, unseen prediction problems at inference time, removing the need for task-specific fine-tuning.
This approach was pioneered by the transformer-based TabPFN [16]. Pretrained on large amounts of synthetically generated classification tasks, its latest incarnation TabPFNv2 [17] produces SOTA results on tabular datasets with up to 10 0001000010\,00010 000 samples for both classification and regression tasks. In recent work, TabICL [28] extends the success story of this approach to even larger datasets, using special tabular embedding modules improving on the quadratic scaling in both the number of features and rows present in the TabPFN architecture.
Common to both TabPFN and TabICL is that they are trained entirely on synthetically generated numerical data, with categorical features produced by an indexing procedure. While using synthetic data has many advantages, in particular its diversity at scale, a consequence is that the data does not contain any semantically meaningful values as found in real-world applications, both in the form of column names, and categorical or free-text entries. Furthermore, such data contains no additional data types such as times or dates that are abundant in practice. Consequently, these models do not utilize such information in a semantically meaningful way even when it is present at inference. In particular, column names are not used in either TabPFN or TabICL, and categorical features are encoded via one-hot or ordinal encoding disregarding any underlying semantics. We argue that semantic understanding can be captured by training on a large number of real-world tabular datasets. A primary example of this philosophy is TabuLa-8B [11], turning the pretrained large-language model (LLM) Llama 3-8B [7] into a tabular ICL model by fine-tuning it on around 3 million tables of the T4 dataset proposed therein. However, utilizing pretrained LLMs for tabular tasks has several limitations: Most importantly, textual serialization and tokenization of the input table is not token efficient, effectively limiting the maximum context length that can be processed. For example, TabuLa-8B operates on a maximum of 32 context rows. Furthermore, the tokenization schema and autoregressive nature of LLMs are not adapted to the tabular structure, resulting in a linear non-uniform token sequence, as cell values can be tokenized into different amounts of tokens, losing the 2D structure of the underlying data. Finally, the serialization and autoregressive processing result in an architecture that is neither row nor column permutation invariant – a property often desirable for tabular data [32].
Aiming to bridge these approaches, we propose a table-native ICL model trained on the real-world T4 dataset [11], using embeddings tailored to different data modalities, in particular incorporating semantic embeddings of column names and categorical values. The resulting model is competitive to existing table-native ICL approaches across a range of tabular prediction benchmarks (OpenML-CC18 [1], OpenML-CTR23 [10], TALENT [37], and TabReD [30]), and achieves a new SOTA for the semantically rich CARTE benchmark [22], in particular in the low-data regime.
2 Related Work
Tabular deep learning:  Prediction on tabular data has traditionally been dominated by decision tree algorithms, particularly boosted variants like XGBoost [2], LightGBM [21], and CatBoost [27]. These models deliver strong performance but require separate training for each dataset and cannot leverage pretraining. Hence, much work has been done to transfer the success of deep learning methods and general pretraining to the tabular setting. Early examples include the FT-Transformer [13] and Xtab [39], whereas only more recent approaches have successfully shown consistently good performance overtaking boosted trees, for example TabR [12], RealMLP [18], or CARTE [22].
In-context learning on tabular data:  TabPFN [16] broke the long-standing dominance of boosted trees on small classification tasks, outperforming them by using row-level ICL. Pretrained on a large amount of synthetic tabular data, it generalizes to new tasks at inference time without fine-tuning or hyperparameter optimization. A recent variant, TabDPT [24], showed that equally excellent results can be achieved by training on real-world data using similarity-based retrieval for the context examples – an idea previously investigated in the TabR approach [12] – and further unlocked regression in this setting. Generalizing the row-based encoding, cell-based ICL introduced with TabPFNv2111In the following, we will focus solely on TabPFNv2 and refer to it simply as TabPFN. [17] and utilized also by TabICL [28] extended this success to larger datasets with up to 10 0001000010\,00010 000 and more samples, even outperforming SOTA AutoML solutions such as AutoGluon [9] on certain benchmarks.
Semantics and real data:  Capturing the rich semantics of real-world tabular datasets is a desirable property of a tabular foundation model, enabling the transfer of world knowledge across prediction tasks in addition to statistical patterns. The CARTE [22] architecture enables pretraining across a variety of real-world sources while capturing table semantics. It achieves SOTA results on the semantically rich CARTE benchmark, although it requires task-specific fine-tuning.
Modern LLMs have both deep semantic understanding and extensive world knowledge. Several works approach tabular ICL by tuning LLMs on tabular tasks, e.g. TabLLM [15], LIFT [6], or TabuLa-8B [11]. In particular, the works by Gardner et al. [11] are note-worthy for curating the T4 dataset, containing roughly 3 M tables extracted and processed from the TabLib collection [8] and for its excellent results in the very low data-regime.
3 Method
Figure 1: Our proposed model architecture illustrating the integration of data type-specific embeddings, an interleaved attention backbone, and customized output heads.
To overcome the aforementioned limitations of existing table-native ICL methods and bridge the gap to LLM-based ones, we propose ConTextTab, a semantics-aware table-native ICL model. To this end, we perform several key modifications to the TabPFN architecture and utilize large-scale pretraining on real-world data. An overview of our proposed architecture is given in Figure 1.
3.1 Encoding
We encode data differently depending on its modality – i.e. text, date, or numeric type. Column headers are also encoded, playing the role of positional encodings as used in TabPFN or TabICL.
Text:  We transform each text cell to an embedding vector using a pretrained text embedding model. Note that we apply this to both free text columns as well as categorical columns, which can then retain the meaning in their labels. Any off-the-shelf embedding model can be used for this purpose – e.g. many are available based on the BERT architecture [5]. As many values have to be embedded for each table, there is a natural trade-off between accuracy and speed. We settle for a comparably small and fast model such as all-MiniLM-L6-v2 [19, 34], since the amount of semantic meaning it captures is already much richer than conventional categorical encodings, and we defer to Section 5.1 for an experimental ablation. Since the embeddings are potentially of a different dimension than the target embedding dimension $d$, we apply a learnable linear layer to the text embeddings.
Date:  In order to endow our model with knowledge of both the relative meaning of a dates (e.g. being able to compare two dates) and special dates (e.g. recurring holidays), we embed each of the numbers representing day, month, and year separately and sum the three resulting vectors. In particular, this is more token efficient then encoding dates into multiple features as is common in established preprocessors such as the one used by AutoGluon.
Numerical:  As numbers do not contain semantic meaning beyond their value, we apply a one-dimensional encoding. To make this procedure more robust during training, first, we clip columns between the 2% and 98% quantiles of the distribution. Second, we scale them to have zero mean and unit variance as is common. By Chebyshev’s inequality, this bounds resulting values to the interval $(-7.1,7.1)$, thus avoiding exploding gradients during training. Finally, the resulting number is multiplied by a learnable vector and a bias is added. If the original value was NaN, 0 is used instead, so the bias works as an “is-NaN” flag. An alternative embedding scheme is described in 3.4.
Column headers:  We embed column headers with the same model used for text cells. The result is passed through a separate learnable linear layer to map to the correct target dimension and summed with the cell embedding. Note that all the above embeddings are fully equivariant under permutations of either rows or columns – a property that is often desirable in the tabular domain. This equivariance makes the predictions more reproducible and robust and eliminates some of the need (such as the mapping of category to id, or of column to id) for bagging in models such as TabPFN. Semantic embeddings of cell values have been previously investigated in other works [35, 36, 22, 31], however, details of how the information is consumed differ in each implementation.
3.2 Backbone
We leave the TabPFN architecture mostly unchanged, with alternating “horizontal” (cross-column) and “vertical” (cross-row) self-attention transformer layers. In cross-column attention, each row is considered as a different batch element, and conversely. Following TabPFN, cross-column attention has no masking, while cross-row attention is masked so that each row can only attend to the provided context. To increase the modularity of our code, the feedforward MLP block of the transformer encoder is repeated after each self-attention, so that “horizontal” and “vertical” blocks have exactly the same weights. For sizing, we use the nomenclature popularized by BERT, e.g. by calling “base” the variant with 12 layers and a hidden dimension of 768. However, the real number of (non-embedding) weights is twice that of a BERT model, due to the presence of both “horizontal” and “vertical” layers.
Model weights can also be optionally shared between each instance of the transformer block, which consists of two interleaved attention layers. Such an architecture can be interpreted as a recurrent neural network (RNN), but unrolled in depth rather than in time. Such an iterative structure results in parameter efficiency and the possibility of stacking more blocks in comparison to the traditional approach. Empirically, we observed that sharing weights did not affect model performance and thus we use weight sharing as the default option for our model in the following.
3.3 Decoding
Classification:  We apply a standard cross-entropy loss after an MLP with at least as large an output dimension as the number of classes. This, however, imposes two limitations. First, the number of classes seen in pretraining cannot be exceeded at inference, without resorting to suboptimal schemas such as hierarchical classification, as used in TabPFN’s many-class extension, TabDPT, or TabICL. Secondly, it prevents us from using the semantic value of the class label. In fact, for the model to know what, say, class “0” means, it needs to build an a priori knowledge of that class ID. Therefore, we must create an additional special input encoding, only for the target column, in addition to the ones described in Section 3.1. Even though this breaks equivariance under permutation, we retain it as we found it to be effective in the most common scenario of few-classes classification.
Regression:  The model predicts the floating-point value of the target, clipped and normalized as described in Section 3.1. Empirically, we have found this simple schema to work well. During training, an $L_{2}$ loss is applied, and during inference, only the inverse transformation of normalization is applied to the prediction.
3.4 Alternative architectures
Number encoding and decoding via binning:  Encoding – soft binning. We split the numbers into bins defined via quantiles to ensure uniform distribution. In order not to lose any fine-grained information while keeping the number of bins limited, we use soft binning: each of the, say $n$, bins is associated with a quantile $\{q_{i}\}_{1,\dots,n}$ of order $\frac{2i-1}{2n}$. Then, any point $x\in[q_{1},q_{n}]$ is encoded via the linear combination $\lambda v_{i}+(1-\lambda)v_{i+1}$, where $x=\lambda q_{i}+(1-\lambda)q_{i+1}$, $0\leq\lambda\leq 1$, and $v_{1},\dots,v_{n}$ are learnable vectors. Anything outside that interval is mapped to either $v_{1}$ or $v_{n}$.
Decoding – bin averaging. The regression is converted to a classification task for the bin it belongs to. During training, cross-entropy loss is applied with hard labels. At prediction time, this provides a probability distribution across all bins, $(p_{1},\dots,p_{n})$; the prediction is then given by $p_{1}q_{1}+\dots+p_{n}q_{n}$.
Supervised clustering head:  We introduce an alternative method to perform classification that has the advantage of retaining semantic meaning not only for features but also for labels. As further benefits, full equivariance is preserved, without the need to map classes to IDs, and the limitation on the number of classes supported for prediction is lifted, improving upon the constraint currently present in many tabular ICL approaches such as TabPFN, TabDPT, or TabICL. For this, we take inspiration from Polewczyk and Spinaci [26]: For each row, the final output corresponding to the target column is mapped by a two-layer MLP to a vector $x$. We then build a matrix of shape $(n_{\mathrm{query}},n_{\mathrm{context}})$ by computing the cosine similarities between these vectors $x_{i}$, $x_{j}$ for each query row $i$ and context row $j$. During training, this is compared against the adjacency matrix with value 1 if two rows belong to the same class, and 0 otherwise: We compute the element-wise binary cross-entropy loss between this adjacency matrix’s entries and the clipped cosine similarities. That is, if $x_{i}$ is the vector from above for row $i$ and $c_{i}$ is its class, then the loss is given by
$\text{Loss}=-\sum_{\begin{subarray}{c}i\in\mathrm{query}\\ j\in\mathrm{context}\end{subarray}}\log(s_{ij})\delta_{c_{i}=c_{j}}+\log(1-s_{ij})\delta_{c_{i}\neq c_{j}},\quad s_{ij}=\text{clip}\Big{(}\frac{x_{i}\cdot x_{j}}{\|x_{i}\|\|x_{j}\|},\varepsilon,1-\varepsilon\Big{)}$ (1)
including a small margin $\varepsilon>0$ for numerical stability. In this way, rows in the same class are pushed to similar embeddings, while rows in different classes can be either opposite or orthogonal. This enables the use of the same semantic-preserving text encoding for the target class of context rows as used for column names and input features, while also supporting an arbitrary number of target classes. A possible use case for this scenario is a situation with many target classes with very few rows per class, allowing extraction of information from semantically similar target classes.
Induced Set Attention Blocks:  One of the main limitations of the standard multi-head attention is its quadratic complexity. While memory can be effectively compressed to linear [29, 4], that is not the case for runtime. This poses a significant challenge when dealing with large tables. Taking inspiration from [23, 20, 28], we experiment with methods to handle larger context more efficiently.
Our proposed architecture replaces some of the multihead attention blocks within TabPFN with ISAB blocks from [23]. Following the notation from [23], let $\mathrm{MAB}(X,Y)$ denote the original multihead attention block from the transformer architecture, followed by a feedforward block including skip connections and layer normalization. Let $I$ be so-called inducing points [23], an additional set of parameters of shape $(n,d)$ for an arbitrary $n$, with $d$ denoting the hidden dimension, and define
$\mathrm{ISAB}(C\mathbin{+\mkern-10.0mu+}Q)=\mathrm{MAB}\big{(}C\mathbin{+\mkern-10.0mu+}Q,\mathrm{MAB}(I\mathbin{+\mkern-10.0mu+}Q,C\mathbin{+\mkern-10.0mu+}Q)\big{)},$ (2)
[…]
Since tables can have a very large number of rows but rarely exceed a few hundred columns, we apply the ISAB block only to cross-row attention. In our architecture, we only use the ISAB block for the first $m$ blocks (e.g., $m=3$), followed by one $\textnormal{MAB}(I\mathbin{+\mkern-10.0mu+}Q,C\mathbin{+\mkern-10.0mu+}Q)$ block, and finally use standard attention for the following layers. This both improves training stability and reduces inference time because the corresponding attention has a bounded number of tokens: $\textnormal{MAB}(I\mathbin{+\mkern-10.0mu+}Q,I\mathbin{+\mkern-10.0mu+}Q).$
4 Experimental Setup
4.1 Training and inference
For pretraining, we use the T4 dataset [11]. We discard tables with fewer than 150 rows, which leaves $2.18$ M tables with a median of 750 rows and 9 columns. We randomly select 1000 rows, then between 50 and 900 rows as query, and use the rest as context. Then we randomly select one target column, excluding all date columns, numerical columns with more than 50% NaN values, and other columns having more than 20% of unique values. Finally, we up-sample non-numeric columns to have roughly the same proportion of regression and classification tasks. We train each model for between 4 and 10 million steps (i.e., 2 to 5 epochs) until convergence. We use a micro batch size of 1 and accumulate gradients to simulate a batch size of 256 (or 128 for smaller models of “mini” size). To improve stability, we employ gradient clipping and the AdamW optimizer with a maximum learning rate of $10^{-4}$, reached after a linear warm-up phase of 1000 gradient updates. Under this setup, we train a “base” sized model on a single H100 GPU, reaching a throughput of roughly 10 tables/s. Hence, full training takes between 4 and 12 days depending on the number of steps. Using the default parameters $n=12$, $d=768$, $d_{ff}=3072$ would result in a total of 172M parameters, but due to weight sharing there are only 16M trainable parameters.
We also experimented with curriculum learning, by further adding in a second step using the same training data as Ma et al. [24]. This data has fewer tables, 123, but many more cells in each table with a median of 11 k rows and 34 columns. In this second step, we increased the number of rows used for training to 4000. We refer to Section 5.1 for an analysis of the impact of training data size on model performance, which we found to be crucial.
At inference time, we apply 8-fold bagging, similar to the ensembling used by TabPFN. That is, from the original train split of a given evaluation dataset, we sample 8 times $c$ context rows with replacement, make a separate prediction with each collection of $c$ rows using the same model, and average the predictions (regression values or classification probabilities). While, during the default training, $c$ never surpasses 950, we can scale this during evaluation as much as memory permits. The combination with bagging allows ConTextTab to use up to $8c$ points as context. To limit runtime, in each bag, we also sample up to 500 columns if the original dataset has more than that.
4.2 Evaluation
Datasets:  We use a variety of tabular prediction datasets to evaluate and compare our approach to established baselines and other SOTA methods. Namely, we evaluate all models on the following benchmarks: OpenML-CC18 [1], a pure classification benchmark; OpenML-CTR23 [10], a pure regression benchmark; TALENT [37], a recently introduced diverse benchmark containing over 300 classification and regression benchmarks. Here, we focus on a subset containing 45 datasets that are representative of the overall performance of the baselines investigated in the original works, which we refer to as the TALENT-Tiny benchmark; TabReD [30], a small but challenging benchmark of large datasets representative of practical prediction tasks; and finally CARTE [22], a mixed classification and regression benchmark containing highly semantic features and few numerical ones.
Across all benchmarks, we evaluate 91 regression and 112 classification tasks, ranging from 5 to 3 k columns and 400 to roughly 400 k training examples. Due to the large number of evaluated datasets, we do not perform cross-validation but evaluate a fixed test split for each task. For models that do not explicitly use a validation split, we concatenate the train and validation splits. We refer to Appendix B.2 for a visualization of the dataset statistics and further details.
Baselines:  We compare our approach to several established classical methods as well as recent ICL and other deep learning models. In particular, we compare with TabPFN [17], TabICL [28], and TabDPT [24] as the most recent table-native ICL methods. For all pretrained models, we use the latest available release and checkpoints as of May 2025. As the SOTA on the CARTE benchmark, we also evaluate the CARTE model proposed by [22]. As another well-performing recent deep learning approach, we evaluate RealMLP [18]. Note that both CARTE and RealMLP need to be finetuned or fitted for each dataset separately. We perform per-dataset hyperparameter tuning via random search for RealMLP as provided in the author’s implementation via the pytabkit package. Also using the implementation wrapper from pytabkit, we evaluate SOTA boosted tree baselines, namely XGBoost [2], LightBGM [21], and CatBoost [27] – both in the meta-tuned default variants (TD) [18] as well as via a Parzen-tree estimator-based hyperparameter optimization (HPO). Furthermore, we evaluate several common, non-tuned baselines from scikit-learn [25], namely naive, linear, and KNN estimators, as well as a random forest and histogram-based gradient boosting tree estimators. Finally, as the gold standard in tabular prediction, we evaluate AutoGluon [9], an AutoML solution stacking and ensembling many of the baselines outlined above. Please refer to Appendix B.1 for full details on the used baseline versions and parameters as well as details about the dataset preprocessing and encoding.
Metrics:  Throughout, we show (mean) accuracy for classification tasks as well as (mean) $R^{2}$ score for regression tasks. As averaging across a large number of datasets with varying performance can blur relative performance across models, we also evaluate (mean) rank. Within each benchmark, we calculate the mean rank across all constituent datasets and models. Full rank performance is averaged across all datasets across all benchmarks, not as per-benchmark averages as some benchmarks contain only few tables. To reduce the impact of small noise, we count model performances as ties when their evaluation scores lie within 0.005 of each other.
5 Results
Table 1: Performance comparison across all evaluated benchmarks, depicting mean accuracy (Acc) for classification and $R^{2}$ score for regression tasks, in percent. Missing values, due to architectural limitations or failed evaluations, are denoted as N/A and excluded from the mean rank calculation.
Model Name All CARTE OML-CC18 OML-CTR23 TabReD TALENT-Tiny Rank Rank Acc R2 Rank Acc Rank R2 Rank Acc R2 Rank Acc R2 AutoGluon 78.7 73.8 88.5 67.0 86.0 64.6 87.9 73.7 ConTextTab 2.20 1.49 76.9 72.2 2.79 86.8 3.29 72.9 3.88 85.4 63.4 2.30 87.7 76.1 TabPFN 2.66 5.08 72.4 65.0 2.67 87.0 2.63 74.9 3.00 85.6 63.8 1.73 87.3 75.1 LightGBM [HPO] 2.85 4.06 72.8 66.1 3.26 86.9 3.31 61.9 1.62 85.9 64.4 3.19 86.4 72.4 XGBoost [HPO] 3.16 4.47 72.6 65.7 4.00 86.3 3.17 71.4 1.75 86.0 64.2 3.24 86.2 72.6 CatBoost [HPO] 3.24 4.02 75.6 66.1 3.57 86.7 4.40 -44.4 2.00 85.8 63.7 3.76 86.0 72.1 HistGradBoost 3.46 5.08 72.5 64.8 3.62 86.1 4.80 51.2 1.75 85.9 63.9 3.62 86.3 67.6 RealMLP [TD] 3.91 8.39 70.2 59.6 3.18 87.2 3.80 45.7 1.25 86.0 64.6 4.08 86.2 71.5 Random forest 4.56 7.08 71.5 63.3 4.69 85.7 5.17 53.6 5.38 85.4 60.7 4.54 85.8 70.6 Naive 8.61 11.22 53.0 -1.8 9.39 47.0 9.69 -8.5 8.62 80.8 -0.6 9.97 53.4 -22.3 CARTE N/A 3.16 76.1 68.5 N/A N/A N/A N/A N/A N/A N/A N/A N/A N/A TabDPT N/A 5.02 72.7 65.1 2.71 87.8 2.46 72.4 N/A N/A 60.9 2.65 86.7 74.8 TabICL N/A N/A 72.5 N/A N/A 88.0 N/A N/A N/A 85.1 N/A N/A 87.4 N/A
The main results, evaluating ConTextTab and the baselines as outlined in the previous section, are shown in Table 1. Here, we exclude AutoGluon from the overall rank and best-model comparison as it is ensembling and stacking a multitude of model architectures, making it difficult to highlight architectural strengths and weaknesses. Overall, our model achieves the highest rank across all evaluated datasets. This performance gain is statistically significant, as depicted in the critical difference diagram in Figure 2 (left). Our model performs particularly well for the semantically rich CARTE benchmark, highlighting the importance of incorporating semantic understanding into tabular ICL. Note that TabPFN, not incorporating semantics, performs worse than both tuned and non-tuned boosted trees in this particular instance. On the other end, while the absolute performance of ConTextTab for large datasets, e.g. within the TabReD benchmark, is not bad, it falls behind tuned boosted trees here, which still achieve SOTA, on par with AutoGluon.
To investigate the dependence on dataset size more closely, we show the performance on the CARTE benchmark across varying subsampled sizes, ranging from 128 rows to the full dataset, in Figure 3. ConTextTab consistently outperforms other models across all sample sizes and even surpasses AutoGluon for up to 2048 training samples. This highlights the strong capabilities of tabular ICL but also the need for further research into scaling these architectures to effectively deal with much larger context sizes as well as training datasets. We investigate the dataset-size dependent rank performance across all 203 datasets in Appendix A.1, further highlighting these challenges.
Additional results, including additional baselines and win ratios, are given in Appendix A.2.
Figure 2: Left: critical difference diagram between ConTextTab and several baselines, across all 203 evaluated datasets. Right: Impact of pretraining dataset size on validation accuracy and $R^{2}$ scores.
Figure 3: Average rank, accuracy, and regression results on the CARTE benchmark across various data subsets, ranging from 128 rows to the full size.
5.1 Ablation studies
In the following, we discuss several experiments ablating aspects of our model reported in Table 2.
Table 2: Multiple ablations as relative differences with respect to one of two base configurations, using binning (top) or one-dimensional embedding (bottom) for regression. Improvements over the base model are highlighted in bold.
Experiment All CARTE OML-CC18 OML-CTR23 TabReD TALENT-Tiny Rank Rank Acc R2 Rank Acc Rank R2 Rank Acc R2 Rank Acc R2 base model (binning) 2.28 1.76 76.0 71.4 2.57 86.8 2.66 70.9 1.00 85.4 62.9 2.32 87.2 73.5 large size 2.76 2.18 +0.2 -0.2 4.01 -5.3 1.77 -0.7 3.75 -25.4 -13.5 1.84 -0.1 -0.0 medium size 5.44 8.45 -5.6 -3.7 4.76 -1.2 3.49 -1.0 5.75 -9.3 -3.5 4.41 -2.6 -2.4 small size 7.68 9.98 -2.9 -6.8 6.57 -3.0 8.17 -4.5 8.12 -1.7 -3.8 6.14 -2.7 -6.7 mini size 8.10 10.69 -4.2 -8.4 7.15 -4.0 8.06 -4.2 5.50 -0.3 -3.5 7.00 -3.7 -6.9 multilingual-e5-small 3.22 4.16 -0.3 -1.5 3.35 -0.4 2.77 -0.1 2.38 -0.2 -1.3 2.27 +0.7 -0.7 gte-multilingual-base 3.22 3.59 +0.0 -1.2 3.18 -0.3 3.60 -1.4 1.62 -0.2 -0.9 2.76 +0.4 -1.3 curriculum learning 2.28 1.51 +1.1 +0.1 3.10 -0.2 1.80 -0.1 1.38 +0.4 -0.2 2.38 -1.0 +0.4 clustering N/A N/A -0.1 N/A 3.82 -1.5 N/A N/A N/A -0.6 N/A N/A -1.3 N/A w/o bagging 2.82 2.76 -0.4 -0.4 3.26 -0.7 3.03 -0.4 1.00 +0.1 -0.5 2.24 -0.6 -0.0 ISAB 5.48 6.16 +0.3 -3.1 5.57 -2.1 5.51 -0.9 3.88 -0.2 -2.7 4.70 -0.5 -6.1 non-shared weights 4.36 5.61 -0.1 -2.6 4.29 -0.9 3.09 +0.8 5.88 -1.0 -2.6 3.65 -0.7 -1.2 base model (1-dim) 2.00 1.80 76.7 72.0 2.36 86.9 2.06 65.0 1.75 85.3 63.3 1.54 87.9 75.8 0.1% clipping 2.03 1.80 0.0 0.0 2.36 0.0 2.37 -0.3 1.38 0.0 +0.2 1.54 0.0 0.0 0.5% clipping 2.08 1.76 0.0 0.0 2.32 -0.1 2.60 +4.0 1.38 +0.1 +0.1 1.70 +0.1 -0.2 2% clipping 2.97 2.88 0.0 -0.5 2.43 -0.1 5.03 +5.8 1.75 +0.1 -0.5 2.43 0.0 -2.0 binning 4.04 4.69 -0.7 -0.6 3.21 -0.1 6.34 +5.9 1.12 +0.1 -0.3 3.24 -0.7 -2.3 context=20000 2.06 1.39 +0.3 +0.6 2.62 -0.2 1.66 +4.2 3.25 -0.4 -0.5 2.00 -0.4 +0.5 context=10000 2.00 1.61 +0.3 +0.3 2.64 -0.2 1.57 +4.2 1.75 0.0 -0.1 1.76 -0.2 +0.5 context=4096 2.81 3.55 -0.1 -0.6 3.03 -0.3 2.11 +4.0 2.25 -0.3 -0.1 2.16 -0.4 +0.4 context=2048 4.11 6.41 -1.0 -1.9 3.61 -0.4 3.06 +3.6 2.50 -0.1 -0.7 3.24 -0.9 +0.1 context=1024 5.49 9.10 -1.3 -3.8 4.33 -0.5 4.60 +2.7 2.62 -0.4 -1.6 4.24 -1.3 -0.6 fine-tuning 2.60 2.51 +0.3 +0.1 3.1 +0.2 1.94 +7.2 2.62 +0.2 -0.3 2.38 -0.1 +0.8
Model size:  As expected, decreasing from “base” to smaller models impacts performance. On the other hand, increasing to “large”, thus doubling the number of layers and increasing the hidden dimension, has no statistically significant impact on performance. As observed in Figure 2 (right), increasing the amount of training data leads to stagnating validation performance gains. We hypothesize that, in the current setup, our model is likely limited by the amount or diversity of data available with T4.
Context size:  The trained models can scale the input context size at inference. As expected, longer context improves results in a monotonic way, in exchange for longer computation time. In practice, the effects are stronger on larger datasets, simply because shorter context models (even with bagging) only get to see a subset of the training data. We stress again that this is entirely extrapolating outside of the pretraining setup, where the largest observed context has size 950. Nevertheless, with larger datasets and context, the model performance saturates, requiring further research in how to increase the capacity of tabular ICL.
Training data:  We ran training on randomly chosen subsets of the T4 dataset of $10^{n}$ tables, $n\in\{1,\dots,6\}$. For this experiment we report the validation scores (on the validation held-out subset of the CARTE dataset) in Figure 2 (right). The results suggest that at least 100 k of the tables are needed in order to train a model of state-of-the-art quality. While exact numbers may vary when using other datasets or model sizes, this gives a high-level estimate of the amount of required data, giving indication of why our model could not profit from a second step of curriculum learning on the much smaller dataset used by Ma et al. [24]. Potentially, training is still limited by data amount or diversity, e.g. missing longer and wider tables, likely required to further scale tabular ICL.
Classification target:  We compare the default cross-entropy loss for classification with the clustering approach described in Section 3.4, allowing to retain the semantics of the target labels, lifting the limitation on the number of target classes as present in TabPFN or TabICL. However, most evaluation datasets only contain non-descriptive labels, and the number of classes is limited. Hence, the potential of clustering cannot emerge, and the default maintains a modest edge. We believe that semantically richer high-cardinality benchmarks are required to spur research in this direction.
ISAB block:  We observed comparable results on classification tasks and a modest drop in performance on regression tasks, while substantially enhancing efficiency by eliminating quadratic time complexity and reducing runtime by a factor of ten for large contexts.
Bagging:  As expected, using bagging at inference time leads to a consistent improvement of results. At the same time, in absolute terms, the gap is rather marginal, probably thanks to the fact that ConTextTab gets rid of most sources of non-equivariance in the architecture (e.g. category indexing and positional encoding). As a result, depending on the use case, this allows to drop bagging, or reduce the number of bags, without compromising performance significantly.
(No) weight sharing:  We observed a slight performance degradation, which may be due to either an insufficient training duration or suboptimal selection of training hyperparameters.
Further ablations:  Other changes (the used sentence embedding model, longer context in training, target clipping for regression, as well as benchmark-dependent fine-tuning) did not result in significant performance gains. Therefore, we refer to the Appendix A.3 for further discussion of these ablations.
6 Conclusions
Limitations and future work:  While achieving SOTA results across the investigated datasets, we observe several limitations of our proposed approach. One drawback of using real-world data for training is the possibility of contamination, e.g. the presence of evaluation tasks in the training corpus. Since the CARTE benchmark is our focus, we conducted a contamination study, using the column name and cell value embeddings created by a text embedder and matching similarities. We did not find any contamination of CARTE in T4. For contamination of OpenML datasets, we refer to the original study by [11]. Either way, as a single table is only seen a few times during training, we believe that memorization is likely not a practical problem when training on real-world data.
Generally, all investigated table-native ICL models, e.g. ours, TabPFN, TabDPT, and TabICL, fail to scale their performance to very large datasets. Increasing the context length did not resolve these issues. Using local context, as done in TabDPT, might overcome such limitations but we were unable to evaluate TabDPT on TabReD as it could not handle large tables. In the large-data case, conventional methods, in particular when stacked and ensembled via AutoGluon, still perform best. Overcoming this remains one of the major research challenges for tabular foundation models.
We observe little gain when training on larger data corpora, e.g. increasing the pretraining subset of T4 or adding tables from the AutoML benchmark. Together with the investigated scaling of our model, we believe that more diverse real-world data is needed to fully unlock semantic understanding at scale. While models trained on synthetic data do not face this issue in principle, how to incorporate semantic alignment in these models remains an open question. Similarly for evaluation, semantically rich benchmarks are scarce. While CARTE includes rich semantic features, all classification targets are binary, containing no semantic information. Furthermore, it contains mostly small to medium sized datasets, in particular in terms of the number of features. Hence, we argue that more diverse, semantically rich data is required – both at scale for pretraining larger models with longer context, as well as curated for evaluation.
Finally, our investigation to utilize target semantics and lift the strict class limit at inference prevalent in current tabular ICL, by using a dynamic supervised clustering approach, did not result in performance gains. This might relate to the lack of semantically meaningful targets in current tabular benchmarks as well as the limited target cardinality across them.
Summary:  We have introduced ConTextTab – a context aware table-native in-context learner trained on real-world data for both classification and regression tasks. Evaluated across a wide range of tabular benchmarks – including both classification and regression tasks – our model achieves state-of-the-art results significantly outperforming existing table-native ICL approaches such as TabPFN, in particular on the semantically rich CARTE benchmark and low-data regime. We outline future research efforts required to further improve semantic understanding in tabular foundation models, in particular the need for more diverse semantically rich pretraining as well as evaluation data.
Acknowledgements
We would like to thank Johannes Hoehne, Johannes Hoffart, and Markus Kohler for their insightful comments and suggestions throughout the development of this work, which have greatly contributed to shaping its direction and quality.
Appendix A Further Results
A.1 Relation between dataset size and model performance
We plot the average rank of each model as a function of the dataset size (expressed in number of rows) across all evaluated datasets in Figure 4. Note that TabICL is missing from this comparison as it only handles classification tasks. We can observe that, as expected, ConTextTab and TabPFN excel in the low data regime, while AutoGluon takes the lead when enough data is available. In the very low data regime below 1000 training rows, TabPFN performs best, however our model surpasses its performance for larger datasets with more than 1000 rows. Overall, ConTextTab remains competitive with AutoGluon until roughly 10 000 rows, possibly as a result of the limitation in both the inference context size and the context size seen during training. After 10 000 rows, gradient boosting methods, as well as RealMLP, start surpassing the performance of ConTextTab as well as TabPFN. Overall, this indicates the need of further research for tabular ICL to handle larger amounts of training data, but also the availability of more diverse benchmarks, covering larger datasets, as the current evaluation is dominated by datasets with less than 10 k rows and less than 100 columns, see Figure 6.
Figure 4: Relation between number of training dataset rows and performance, obtained as a LOWESS regression in the plane $\log(n_{\mathrm{rows}},\textnormal{rank})$. The confidence bands are the 80% confidence intervals obtained via bootstrapping.
A.2 Extended results
Including further baselines not depicted in the main paper, Table 3 shows all collected evaluations across the investigated benchmarks. Furthermore, we show the win ratio confusion matrix and averages of the investigated models in Figure 5.
(a) Win ratio matrix with Model A wins over Model B.
(b) Average Win ratio of Model A against all others.
Figure 5: Win ratio confusion matrix and average of the investigated models across all 203 datasets. Wins are calculated based on accuracy on classification and $R^{2}$ on regression datasets. Ties are not counted as wins. Models are sorted by descending overall rank. Note that CARTE, TabDPT, RealMLP [HPO], and TabICL were not successfully evaluated on all datasets and only compared on evaluated ones, resulting in a smaller support and skewing their win rates. Therefore, they are separately grouped last, not used in averaging, and only shown for completeness.
Table 3: Performance comparison across all evaluated benchmarks, depicting mean accuracy (Acc) for classification and $R^{2}$ score for regression tasks, in percent. Missing values, due to architectural limitations or failed evaluations, are denotes as N/A and excluded from the mean rank calculation.
Model Name All CARTE OML-CC18 OML-CTR23 TabReD TALENT-Tiny Rank Rank Acc R2 Rank Acc Rank R2 Rank Acc R2 Rank Acc R2 Autogluon 2.48 1.55 78.7 73.8 2.61 88.5 4.06 67.0 1.75 86.0 64.6 2.16 87.9 73.7 ConTextTab 3.53 2.24 76.9 72.2 4.08 86.8 4.37 72.9 4.62 85.4 63.4 3.22 87.7 76.1 TabPFN 4.27 7.25 72.4 65.0 3.61 87.0 3.43 74.9 4.5 85.6 63.8 2.16 87.3 75.1 LightGBM [HPO] 4.59 5.41 72.8 66.1 4.56 86.9 4.63 61.9 1.75 85.9 64.4 4.11 86.4 72.4 CatBoost [HPO] 4.99 5.33 75.6 66.1 4.6 86.7 6.63 -44.4 2.12 85.8 63.7 4.32 86.0 72.1 CatBoost [TD] 4.99 6.35 75.1 65.6 4.96 86.4 4.51 67.2 1.25 85.9 64.3 4.43 85.9 74.8 LightGBM [TD] 5.05 7.02 73.1 64.9 4.01 86.8 6.29 39.7 2.25 85.7 63.2 3.81 86.3 72.6 XGBoost [HPO] 5.05 5.84 72.6 65.7 5.74 86.3 4.34 71.4 1.75 86.0 64.2 4.03 86.2 72.6 XGBoost [TD] 5.52 8.24 72.3 64.4 4.03 87.0 6.54 63.7 2.5 85.6 62.8 4.35 86.0 72.7 HistGradBoost 5.81 7.04 72.5 64.8 5.11 86.1 7.14 51.2 2.25 85.9 63.9 5.0 86.3 67.6 RealMLP [TD] 6.52 12.2 70.2 59.6 4.62 87.2 5.14 45.7 1.12 86.0 64.6 4.86 86.2 71.5 Random forest 7.59 9.82 71.5 63.3 6.57 85.7 7.77 53.6 7.25 85.4 60.7 6.38 85.8 70.6 Linear 13.46 16.55 62.7 19.1 11.17 80.9 14.2 43.8 12.12 80.8 -4269.6 13.27 80.5 41.3 KNN [k=5] 13.78 16.25 65.5 34.3 12.06 81.7 13.57 14.1 13.5 78.7 -15.4 14.0 80.3 60.0 Naive 16.35 18.04 53.0 -1.8 15.28 47.0 16.4 -8.5 12.62 80.8 -0.6 16.89 53.4 -22.3 CARTE N/A 4.65 76.1 68.5 N/A N/A N/A N/A N/A N/A N/A 7.38 84.4 71.1 TabDPT N/A 6.63 72.7 65.1 3.74 87.8 3.26 72.4 N/A N/A 60.9 3.73 86.7 74.8 RealMLP [HPO] N/A 11.78 70.4 60.9 4.08 87.3 4.51 53.2 N/A N/A N/A 4.0 86.3 72.1 TabICL N/A N/A 72.5 N/A N/A 88.0 N/A N/A N/A 85.1 N/A N/A 87.4 N/A
A.3 Further ablations
In the following, we discuss in more detail the ablations as presented in Table 2.
Text embedding model:  We test the performance of two more recently released sentence embedding models, namely Multilingual e5 small [33] and gte-multilingual-base [38]. However, we cannot observe significant impact on the final results over the default all-MiniLM-L6-v2 sentence embedder. This potentially indicates the necessity of larger, semantically rich training data or further research into aligning these sentence embedders with a tabular foundation model.
Curriculum training with longer context:  In this setup, after training the model on T4 with a context size of up to 950, training is extended by including also the dataset used by Ma et al. [24], collected from the OpenML AutoML benchmark, and extending the context size up to 4000 rows. Data is sampled from either T4 with 80% probability or the dataset from by Ma et al. [24] with 20% probability. The resulting model shows some improvements on a few tasks, but by statistically insignificant amounts.
Regression target and clipping:  When using a one-dimensional encoding for numerical values, we can choose to clip the tails of the distribution for better stability. This is needed in pretraining due to the noisy data. In practice, whether to do this on downstream tasks can be chosen depending on outlier analysis of the data itself. We test alternative levels of clipping: no clipping (default), clipping at 0.1%, 0.5% and 2% quantiles. We also compare this to using binning instead of one-dimensional embeddings. We find that both clipping by 2% and binning cause a slight reduction in quality (since we are using $n=16$ bins, note that binning also incurs in clipping at $\frac{1}{32}\approx 3\%$). The only exception is in the OpenML-CTR23 benchmark: it contains one dataset with extreme outliers where no model achieves a positive $R^{2}$ score, and adding clipping there limits degenerate scores.
Impact of fine-tuning:  The training process can be easily adapted to fine-tuning on a specific collection of datasets or even a single task.We experimented with fine-tuning on (the training partition of) all the 203 datasets at the same time.However, we observe that, if during this fine-tuning we fix the target column to always be the one corresponding to the prediction task to be evaluated, we rapidly encountered overfitting and prediction scores degrade. This might be a similar problem to the scaling law of training data, as also observed by Ma et al. [24]. In the present situation, we can avoid overfitting by selecting the target column randomly with the same selection procedure described in Section 4.1. However, as shown in Table 2, the gain is rather limited.
Appendix B Experimental Setup – Details
B.1 Baselines
Data preprocessor:  Evaluating models across a multitude of datasets can be tricky. Datasets may have inconsistent data type annotations, such as categories represented as strings or categorical data types, covering low- and high-cardinality categorical features, date, time, or datetime instances, free text, boolean values and more. Most models, however, require numerical input to process and handle non-numerical data types or missing values differently. To unify our evaluation, we implemented a configurable default feature encoder built on the AutoMLPipelineFeatureGenerator from AutoGluon [9] which we found to be very versatile and robust. In particular, the encoder natively handles low- and high-cardinality categorical data, free text (to some extent), as well as datetime encoding. For flexibility and compatibility with a multitude of models, we extended the default implementation to cover the following options that can be adapted to the capabilities of the baseline model at hand:
  • Whether to convert booleans to string/categorical values
  • Whether to not encode string/categorical values for models that natively handle them, such as CatBoost
  • Whether to scale numerical data via quantile scaling with a normal distribution as target
  • Whether to drop constant features
  • Whether to impute missing values, extending the standard imputation (using most frequent categories and mean for numerical data) to bools and datetime data types
As deault, we choose to convert booleans, encode categoricals via ordinal encoding, scale numerical data, drop constant columns and impute missing values.
Below, we describe for which baselines the default values are changed or when other types of feature encodings are used.
TabPFN:  We use the model from the official Python tabpfn package with version 2.0.8 together with the tabpfn-extensions package version 0.1.0 at commit d44606e35f89e18b6bc4c4a2eef2f46918c4302e of the Git repository [^2] https://github.com/PriorLabs/tabpfn-extensions.git as the PyPi release is not up-to-date.
Naturally, we use TabPFNClassifier and TabPFNRegressor for classification and regression tasks, respectively, using default parameters for both. In particular, TabPFNClassifier uses an ensemble of 4 and TabPFNRegressor an ensemble of 8 estimators. We combine the classification estimator with the ManyClassClassifier extension with a redundancy factor of 4 to enable classification beyond the native 10-class limit of TabPFN which is required for the evaluation of some of the 203 evaluated datasets.
For datasets larger than the native 10 k limit of TabPFN, we sample a random 10 k subset of the training split. This affects 66 out of the 203 evaluated datasets. For datasets with more than the 500 feature limit that TabPFN was trained with, we select a random subsample of 500 features. This affects 12 out of 203 evaluated datasets. While this is not optimal, and post-hoc ensembling as well as a random forest preprocessing is recommended by the authors [17], these extensions cannot be combined with the many-class extension required to predict beyond the 10-class limit of the native TabPFN model. Hence, we cannot evaluate TabPFN with the post-hoc ensembling or random forest extension.
As we found the native feature encoder of TabPFN to not work across all evaluated datasets, we use our standard feature encoder (see above), encoding categorical columns, scaling numerical values, dropping constant columns, and imputing missing values. As this procedure should be very similar to the TabPFN-native encoder, we anticipate this deviation to affect the results only insignificantly if at all.
TabICL:  We use the latest model weights tabicl-classifier-v1.1-0506.ckpt from the recent 0.1.2 version of the official tabicl package. This updated variant is an improved checkpoint over the one reported in the original works [28]. As TabICL solely supports classification tasks, we exclude it from the overall mean rank evaluation. For encoding, we use our default encoder, but do not scale numericals, do not drop constant values, and do not impute missing ones as it is natively handled by the model.
TabDPT:  We use the model from the official GitHub repository [^3] https://github.com/layer6ai-labs/TabDPT.git at the most recent 1.1.0 release and tabdpt1_1.pth model checkpoint. Naturally, we use TabDPTClassifier and TabDPTRegressor for classification and regression tasks, respectively, using default parameters for both. Throughout, we evaluate the model with a (local) context size of 2048 which is the best-performing one in the original works [24]. However, evaluation failed for some datasets due to an error in the original code leading to empty predictions for very large datasets in the TabReD benchmark. Here, we use our default encoder, scaling numericals, dropping constant values, and imputing missing ones.
CARTE:  We use the model provided in the official Python carte-ai package with version 0.0.26. We use CARTEClassifier and CARTERegressor with default parameters for classification and regression tasks, respectively We treat binary classification tasks as 2-class multi-class classification and hence set loss=‘‘categorical_crossentropy’’ for the classifacation estimator. With CARTE, we use our default preprocessor to convert bool values and datetime instances and to impute missing values, but otherwise rely on the Table2GraphTransformer provided in the reference implementation.
Pytabkit models:  We use the pytabkit implementation wrapper for evaluating RealMLP, XGBoost, LightGBM, and CatBoost. We evaluate all models both in the tuned-defaults (TD) variant proposed by Holzmüller et al. [18] as well as hyperparameter-optimized (HPO). However, the HPO version did run into OOM issues for RealMLP, even when running on a H100 with 96 GB of VRAM which is why we display the TD variant by default.
In particular, for RealMLP (TD), we use RealMLP_TD_Classifier and RealMLP_TD_Regressor for classification and regression tasks, respectively. For RealMLP (HPO), we use RealMLP_HPO_Classifier and RealMLP_HPO_Regressor for classification and regression tasks, respectively, conducting the default 50 rounds of random search HPO.
For XGBoost (TD), LightGBM (TD), and CatBoost (TD), we use XGB_TD_Classifier, XGB_TD_Regressor, LGBM_TD_Classifier, LGBM_TD_Regressor, CatBoost_TD_Classifier, and CatBoost_TD_Regressor for classification and regression tasks, respectively. For the HPO-variants, we use the HPO_TPE versions of the estimators, performing Parzen-tree based HPO with 50 rounds using the search space as defined by [14]. LightGBM and XGBoost are evaluated on 8-core CPU machines with 64 GB of RAM, whereas CatBoost and RealMLP are evaluated on H100 GPUs with 96 GB of VRAM. Note that CatBoost evaluation on CPU was too slow to evaluate at scale, in particular in the HPO variant. However, there are known issues with the GPU implementation of CatBoost which might degrade performance [^4] See https://github.com/catboost/catboost/issues/1408.. We were not able to observe systematically worse results on those datasets on which we were also able to evaluate the CPU variant. Hence, for consistency, we present results for the GPU variant throughout.
Throughout, we use our default encoder, scaling numericals, dropping constant values, and imputing missing ones. For all models but CatBoost, we perform ordinal encoding of categoricals.
Sklearn models:  We use several standard baseline models from scikit-learn [25], combining them with the default preprocessor as outlined above. Across all scikit-learn baselines, preprocessing only differs in missing value imputation, depending on the model’s capability to handle missing values natively. Throughout, evaluation is performed using scikit-learn v1.5.2.
For the naive predictor, we use the DummyClassifier and DummyRegressor to predict the most frequent, respectively mean value of the train splits as the naive majority baseline.
For the linear predictor, we use the LogisticRegression and LinearRegression for classification and regression tasks, respectively, using default hyperparameters.
For the KNN predictor, we use the KNeighborsClassifier and KNeighborsRegressor for classification and regression tasks, respectively, using default hyperparameters and $k=5$ nearest neighbors.
For the random forest predictor, we use the RandomForestClassifier and RandomForestRegressor for classification and regression tasks, respectively, using default hyperparameters. The model handles missing values natively.
Finally, for the histogram-based gradient boosted tree predictor, we use the HistGradientBoostingClassifier and HistGradientBoostingRegressor for classification and regression tasks, respectively, using default hyperparameters. The model handles missing values natively.
AutoGluon:  Throughout, we use AutoGluon v1.2 and its TabularPredictor without custom preprocessing. We use the best_quality preset and set a per-dataset time limit of 4 h. Otherwise, parameters are left at their default values. For all datasets, evaluation is executed on a single 16-core machine with 128 GB of RAM and no GPU.
B.2 Datasets
Full details of all used dataset and benchmarks are provided in Table LABEL:tab:all-dataset-tables. The row and column count statistics are further visualized in Figure 6.
We extracted all datasets from their original source and performed a custom stratified train-validation-test split with a 70-10-20 ratio. For classification tasks, the target column is used for stratification. For regression tasks, we perform stratification on the binned target column, binning it into 5 quantiles using the qcut method from the pandas library. Otherwise, we do not perform any alterations on the data. Models not using a specific validation procedure are provided with the concatenated train and validation split for training.
Figure 6: Column and row distribution of the evaluated benchmark datasets.
Table 4: Details of all used benchmark datasets.
Benchmark Table Num. rows Num. cols Target type
carte anime_planet 11512 11 regression
carte babies_r_us 4068 5 regression
carte beer_ratings 2557 20 regression
carte bikedekho 3828 8 regression
carte bikewale 7193 8 regression
carte buy_buy_baby 8574 5 regression
carte cardekho 30250 17 regression
carte chocolate_bar_ratings 2070 9 classification
carte clear_corpus 3779 27 regression
carte coffee_ratings 1661 9 classification
carte company_employees 8752 8 regression
carte employee_remuneration 28316 5 regression
carte employee_salaries 7368 9 regression
carte fifa22_players 14468 19 regression
carte filmtv_movies 32964 10 regression
carte journal_jcr 7692 10 regression
carte journal_sjr 22344 10 regression
carte jp_anime 12428 17 regression
carte k_drama 991 13 regression
carte michelin 5465 8 classification
carte mlds_salaries 8364 9 regression
carte movies 5779 14 regression
carte museums 9173 17 regression
carte mydramalist 2720 14 regression
carte nba_draft 1335 7 classification
carte prescription_drugs 1371 8 regression
carte ramen_ratings 3267 5 classification
carte roger_ebert 2149 8 classification
carte rotten_tomatoes 5726 14 regression
carte spotify 32879 19 classification
carte us_accidents_counts 18098 7 regression
carte us_accidents_severity 17324 11 classification
carte us_presidential 15885 7 regression
carte used_cars_24 4734 9 regression
carte used_cars_benz_italy 13112 8 regression
carte used_cars_dot_com 3207 11 regression
carte used_cars_pakistan 58124 8 regression
carte used_cars_saudi_arabia 4405 11 regression
carte videogame_sales 13128 6 regression
carte whisky 1449 7 classification
carte wikiliq_beer 10768 10 regression
carte wikiliq_spirit 9820 8 regression
carte wina_pl 1797 16 regression
carte wine_dot_com_prices 12203 10 regression
carte wine_dot_com_ratings 3276 10 regression
carte wine_enthusiasts_prices 96780 10 regression
carte wine_enthusiasts_ratings 103976 10 regression
carte wine_vivino_price 11067 8 regression
carte wine_vivino_rating 11067 9 regression
carte yelp 51692 13 classification
carte zomato 49152 9 classification
openml_cc18 adult 39073 15 classification
openml_cc18 analcatdata_authorship 672 71 classification
openml_cc18 analcatdata_dmft 637 5 classification
openml_cc18 balance_scale 500 5 classification
openml_cc18 bank_marketing 36168 17 classification
openml_cc18 banknote_authentication 1097 5 classification
openml_cc18 bioresponse 3000 1777 classification
openml_cc18 blood_transfusion_service_center 598 5 classification
openml_cc18 breast_w 559 10 classification
openml_cc18 car 1382 7 classification
openml_cc18 churn 4000 21 classification
openml_cc18 cifar_10 48000 3073 classification
openml_cc18 climate_model_simulation_crashes 432 19 classification
openml_cc18 cmc 1178 10 classification
openml_cc18 cnae_9 864 857 classification
openml_cc18 connect_4 54045 43 classification
openml_cc18 credit_approval 552 16 classification
openml_cc18 credit_g 800 21 classification
openml_cc18 cylinder_bands 432 38 classification
openml_cc18 devnagari_script 73600 1025 classification
openml_cc18 diabetes 614 9 classification
openml_cc18 dna 2548 181 classification
openml_cc18 dresses_sales 400 13 classification
openml_cc18 electricity 36249 9 classification
openml_cc18 eucalyptus 588 20 classification
openml_cc18 fashion_mnist 56000 785 classification
openml_cc18 first_order_theorem_proving 4894 52 classification
openml_cc18 gesturephasesegmentationprocessed 7898 33 classification
openml_cc18 har 8239 562 classification
openml_cc18 ilpd 466 11 classification
openml_cc18 internet_advertisements 2623 1559 classification
openml_cc18 isolet 6237 618 classification
openml_cc18 jm1 8708 22 classification
openml_cc18 jungle_chess_2pcs_raw_endgame_complete 35855 7 classification
openml_cc18 kc1 1687 22 classification
openml_cc18 kc2 417 22 classification
openml_cc18 kr_vs_kp 2556 37 classification
openml_cc18 letter 16000 17 classification
openml_cc18 madelon 2080 501 classification
openml_cc18 mfeat_factors 1600 217 classification
openml_cc18 mfeat_fourier 1600 77 classification
openml_cc18 mfeat_karhunen 1600 65 classification
openml_cc18 mfeat_morphological 1600 7 classification
openml_cc18 mfeat_pixel 1600 241 classification
openml_cc18 mfeat_zernike 1600 48 classification
openml_cc18 miceprotein 864 78 classification
openml_cc18 mnist_784 56000 785 classification
openml_cc18 nomao 27572 119 classification
openml_cc18 numerai28_6 77056 22 classification
openml_cc18 optdigits 4496 65 classification
openml_cc18 ozone_level_8hr 2027 73 classification
openml_cc18 pc1 887 22 classification
openml_cc18 pc3 1250 38 classification
openml_cc18 pc4 1166 38 classification
openml_cc18 pendigits 8793 17 classification
openml_cc18 phishingwebsites 8844 31 classification
openml_cc18 phoneme 4323 6 classification
openml_cc18 qsar_biodeg 844 42 classification
openml_cc18 satimage 5144 37 classification
openml_cc18 segment 1848 17 classification
openml_cc18 semeion 1274 257 classification
openml_cc18 sick 3017 30 classification
openml_cc18 spambase 3680 58 classification
openml_cc18 splice 2552 61 classification
openml_cc18 steel_plates_fault 1552 28 classification
openml_cc18 texture 4400 41 classification
openml_cc18 tic_tac_toe 766 10 classification
openml_cc18 vehicle 676 19 classification
openml_cc18 vowel 792 13 classification
openml_cc18 wall_robot_navigation 4364 25 classification
openml_cc18 wdbc 455 31 classification
openml_cc18 wilt 3871 6 classification
openml_ctr23 abalone 3341 9 regression
openml_ctr23 airfoil_self_noise 1202 6 regression
openml_ctr23 auction_verification 1634 8 regression
openml_ctr23 brazilian_houses 8553 10 regression
openml_ctr23 california_housing 16512 9 regression
openml_ctr23 cars 643 18 regression
openml_ctr23 concrete_compressive_strength 824 9 regression
openml_ctr23 cps88wages 22524 7 regression
openml_ctr23 cpu_activity 6553 22 regression
openml_ctr23 diamonds 43152 10 regression
openml_ctr23 energy_efficiency 614 9 regression
openml_ctr23 fifa 15342 29 regression
openml_ctr23 forest_fires 413 13 regression
openml_ctr23 fps_benchmark 19699 44 regression
openml_ctr23 geographical_origin_of_music 847 117 regression
openml_ctr23 grid_stability 8000 13 regression
openml_ctr23 health_insurance 17817 12 regression
openml_ctr23 kin8nm 6553 9 regression
openml_ctr23 kings_county 17290 22 regression
openml_ctr23 miami_housing 11145 16 regression
openml_ctr23 moneyball 985 15 regression
openml_ctr23 naval_propulsion_plant 9547 15 regression
openml_ctr23 physiochemical_protein 36584 10 regression
openml_ctr23 pumadyn32nh 6553 33 regression
openml_ctr23 qsar_fish_toxicity 726 7 regression
openml_ctr23 red_wine 1279 12 regression
openml_ctr23 sarcos 39146 22 regression
openml_ctr23 socmob 924 6 regression
openml_ctr23 solar_flare 852 11 regression
openml_ctr23 space_ga 2485 7 regression
openml_ctr23 student_performance_por 519 31 regression
openml_ctr23 superconductivity 17010 82 regression
openml_ctr23 video_transcoding 55027 19 regression
openml_ctr23 wave_energy 57600 49 regression
openml_ctr23 white_wine 3918 12 regression
tabred acquire_valued_shoppers_challenge 133602 117 classification
tabred cooking_time 278338 196 regression
tabred delivery_eta 313589 227 regression
tabred home_credit_credit_risk_model_stability 325663 701 classification
tabred homesite_quote_conversion 244458 302 classification
tabred maps_routing 219994 991 regression
tabred sberbank_russian_housing_market 23674 396 regression
tabred tabred_weather 382955 111 regression
talent_tiny ailerons 11000 41 regression
talent_tiny breast_w 31492 10 classification
talent_tiny cmc 44236 10 classification
talent_tiny dis 3017 30 classification
talent_tiny eye_movements_bin 6086 21 classification
talent_tiny fried 32614 11 regression
talent_tiny house_16h_reg 18227 17 regression
talent_tiny ibm_employee_performance 1176 31 classification
talent_tiny jungle_chess_2pcs_raw_endgame_complete 35855 7 classification
talent_tiny kaggle_bike_sharing_demand_challange 8708 10 regression
talent_tiny kc1 1687 22 classification
talent_tiny kin8nm 6553 9 regression
talent_tiny law_school_admission_bianry 16640 12 regression
talent_tiny mfeat_fourier 1600 77 classification
talent_tiny mv 32614 11 regression
talent_tiny okcupid_stem 21341 14 classification
talent_tiny online_shoppers 9864 15 classification
talent_tiny optdigits 4496 65 classification
talent_tiny page_blocks 4378 11 classification
talent_tiny pc3 1250 38 classification
talent_tiny pendigits 8793 17 classification
talent_tiny pol 8065 27 classification
talent_tiny pol_reg 12000 49 regression
talent_tiny rl 3976 13 classification
talent_tiny satimage 5144 37 classification
talent_tiny segment 1848 18 classification
talent_tiny socmob 924 6 regression
talent_tiny splice 2552 61 classification
talent_tiny sylvine 4099 21 classification
talent_tiny thyroid_dis 2240 27 classification
talent_tiny tic_tac_toe 31492 10 classification
talent_tiny vulnonevul 4553 17 regression
talent_tiny waterstress 950 23 regression
talent_tiny waveform_5000 4000 41 classification
talent_tiny website_phishing 1082 10 classification
talent_tiny wine 2043 5 classification
talent_tiny wine_quality_white 3918 12 classification

### v4

@@
 Tabular in-context learning (ICL) has recently achieved state-of-the-art (SOTA) performance on several tabular prediction tasks.
-Previously restricted to classification problems on small tables, recent advances such as TabPFN [17] and TabICL [28] have extended its use to larger datasets.
+Previously restricted to classification problems on small tables, recent advances such as TabPFN [tabpfnv2] and TabICL [tabicl] have extended its use to larger datasets.
 Although current table-native ICL architectures are architecturally efficient and well-adapted to tabular data structures, their exclusive training on synthetic data limits their ability to fully leverage the rich semantics and world knowledge contained in real-world tabular data.
-At the other end of the spectrum, tabular ICL models based on pretrained large language models such as TabuLa-8B [11] integrate deep semantic understanding and world knowledge but are only able to make use of a small amount of context due to inherent architectural limitations.
+At the other end of the spectrum, tabular ICL models based on pretrained large language models such as TabuLa-8B [tabula8b] integrate deep semantic understanding and world knowledge but are only able to make use of a small amount of context due to inherent architectural limitations.
 With the aim to combine the best of both these worlds, we introduce ConTextTab, integrating semantic understanding and alignment into a table-native ICL framework.
 By employing specialized embeddings for different data modalities and by training on large-scale real-world tabular data, our model is competitive with SOTA across a broad set of benchmarks while setting a new standard on the semantically rich CARTE benchmark.
-Code and checkpoints are available at: https://github.com/SAP-samples/contexttab.
+Code and model checkpoints are available at: https://github.com/SAP-samples/sap-rpt-1-oss.
+**footnotetext: Equal contribution.
 1 Introduction
-Tables with information spread across rows and columns remain a predominant data format in many real-world applications [3], making their understanding through machine learning algorithms critical.
+Tables with information spread across rows and columns remain a predominant data format in many real-world applications [chui2018notes], making their understanding through machine learning algorithms critical.
 Despite the great success of deep learning approaches in natural language processing and computer vision, leveraging large amounts of pretraining data, conventional machine learning methods such as gradient boosted trees and their variants remain the predominant state-of-the-art (SOTA) across tabular prediction benchmarks.
-Recently, however, applying the in-context learning (ICL) paradigm to tabular prediction tasks has shown promising results by enabling pretraining of deep learning models across a large set of heterogeneous tables, constituting a new SOTA on small to medium tabular prediction tasks [17].
+Recently, however, applying the in-context learning (ICL) paradigm to tabular prediction tasks has shown promising results by enabling pretraining of deep learning models across a large set of heterogeneous tables, constituting a new SOTA on small to medium tabular prediction tasks [tabpfnv2].
 In this setting, predicting a target value $y$ based on the input features $\boldsymbol{x}$ of a row in a table $T$ additionally uses further rows from $T$ (the context), including their target values.
 This enables the model to adapt to new, unseen prediction problems at inference time, removing the need for task-specific fine-tuning.
-This approach was pioneered by the transformer-based TabPFN [16].
-Pretrained on large amounts of synthetically generated classification tasks, its latest incarnation TabPFNv2 [17] produces SOTA results on tabular datasets with up to 10 0001000010\,00010 000 samples for both classification and regression tasks.
-In recent work, TabICL [28] extends the success story of this approach to even larger datasets, using special tabular embedding modules improving on the quadratic scaling in both the number of features and rows present in the TabPFN architecture.
+This approach was pioneered by the transformer-based TabPFN [tabpfnv1].
+Pretrained on large amounts of synthetically generated classification tasks, its latest incarnation TabPFNv2 [tabpfnv2] produces SOTA results on tabular datasets with up to 10 00010\,000 samples for both classification and regression tasks.
+In recent work, TabICL [tabicl] extends the success story of this approach to even larger datasets, using special tabular embedding modules improving on the quadratic scaling in both the number of features and rows present in the TabPFN architecture.
 Common to both TabPFN and TabICL is that they are trained entirely on synthetically generated numerical data, with categorical features produced by an indexing procedure.
@@
 We argue that semantic understanding can be captured by training on a large number of real-world tabular datasets.
-A primary example of this philosophy is TabuLa-8B [11], turning the pretrained large-language model (LLM) Llama 3-8B [7] into a tabular ICL model by fine-tuning it on around 3 million tables of the T4 dataset proposed therein.
+A primary example of this philosophy is TabuLa-8B [tabula8b], turning the pretrained large-language model (LLM) Llama 3-8B [dubey2024llamareducedauthors] into a tabular ICL model by fine-tuning it on around 3 million tables of the T4 dataset proposed therein.
 However, utilizing pretrained LLMs for tabular tasks has several limitations: Most importantly, textual serialization and tokenization of the input table is not token efficient, effectively limiting the maximum context length that can be processed.
@@
 Furthermore, the tokenization schema and autoregressive nature of LLMs are not adapted to the tabular structure, resulting in a linear non-uniform token sequence, as cell values can be tokenized into different amounts of tokens, losing the 2D structure of the underlying data.
-Finally, the serialization and autoregressive processing result in an architecture that is neither row nor column permutation invariant – a property often desirable for tabular data [32].
-Aiming to bridge these approaches, we propose a table-native ICL model trained on the real-world T4 dataset [11], using embeddings tailored to different data modalities, in particular incorporating semantic embeddings of column names and categorical values.
-The resulting model is competitive to existing table-native ICL approaches across a range of tabular prediction benchmarks (OpenML-CC18 [1], OpenML-CTR23 [10], TALENT [37], and TabReD [30]), and achieves a new SOTA for the semantically rich CARTE benchmark [22], in particular in the low-data regime.
+Finally, the serialization and autoregressive processing result in an architecture that is neither row nor column permutation invariant – a property often desirable for tabular data [towardsfoundation].
+Aiming to bridge these approaches, we propose a table-native ICL model trained on the real-world T4 dataset [tabula8b], using embeddings tailored to different data modalities, in particular incorporating semantic embeddings of column names and categorical values.
+The resulting model is competitive to existing table-native ICL approaches across a range of tabular prediction benchmarks (OpenML-CC18 [openmlcc18], OpenML-CTR23 [openmlctr23], TALENT [talent], and TabReD [tabred]), and achieves a new SOTA for the semantically rich CARTE benchmark [CARTE], in particular in the low-data regime.
 2 Related Work
-Tabular deep learning:  Prediction on tabular data has traditionally been dominated by decision tree algorithms, particularly boosted variants like XGBoost [2], LightGBM [21], and CatBoost [27].
+Tabular deep learning:  Prediction on tabular data has traditionally been dominated by decision tree algorithms, particularly boosted variants like XGBoost [xgboost], LightGBM [lightgbm], and CatBoost [catboost].
 These models deliver strong performance but require separate training for each dataset and cannot leverage pretraining.
 Hence, much work has been done to transfer the success of deep learning methods and general pretraining to the tabular setting.
-Early examples include the FT-Transformer [13] and Xtab [39], whereas only more recent approaches have successfully shown consistently good performance overtaking boosted trees, for example TabR [12], RealMLP [18], or CARTE [22].
-In-context learning on tabular data:  TabPFN [16] broke the long-standing dominance of boosted trees on small classification tasks, outperforming them by using row-level ICL.
+Early examples include the FT-Transformer [ft-transformer] and Xtab [xtab], whereas more recent approaches have shown consistently good performance overtaking boosted trees, for example TabR [tabr], RealMLP [realmlp], CARTE [CARTE], TabM [tabm], or ModernNCA [modernnca].
+In-context learning on tabular data:  TabPFN [tabpfnv1] broke the long-standing dominance of boosted trees on small classification tasks, outperforming them by using row-level ICL.
 Pretrained on a large amount of synthetic tabular data, it generalizes to new tasks at inference time without fine-tuning or hyperparameter optimization.
-A recent variant, TabDPT [24], showed that equally excellent results can be achieved by training on real-world data using similarity-based retrieval for the context examples – an idea previously investigated in the TabR approach [12] – and further unlocked regression in this setting.
+A recent variant, TabDPT [tabdpt], showed that equally excellent results can be achieved by training on real-world data using similarity-based retrieval for the context examples – an idea previously investigated in the TabR approach [tabr] – and further unlocked regression in this setting.
 Generalizing the row-based encoding, cell-based ICL introduced with TabPFNv2111In the following, we will focus solely on TabPFNv2 and refer to it simply as TabPFN.
-[17] and utilized also by TabICL [28] extended this success to larger datasets with up to 10 0001000010\,00010 000 and more samples, even outperforming SOTA AutoML solutions such as AutoGluon [9] on certain benchmarks.
+[tabpfnv2] and utilized also by TabICL [tabicl] extended this success to larger datasets with up to 10 00010\,000 and more samples, even outperforming SOTA AutoML solutions such as AutoGluon [autogluon] on certain benchmarks.
 Semantics and real data:  Capturing the rich semantics of real-world tabular datasets is a desirable property of a tabular foundation model, enabling the transfer of world knowledge across prediction tasks in addition to statistical patterns.
-The CARTE [22] architecture enables pretraining across a variety of real-world sources while capturing table semantics.
+The CARTE [CARTE] architecture enables pretraining across a variety of real-world sources while capturing table semantics.
 It achieves SOTA results on the semantically rich CARTE benchmark, although it requires task-specific fine-tuning.
@@
 Several works approach tabular ICL by tuning LLMs on tabular tasks, e.g.
-TabLLM [15], LIFT [6], or TabuLa-8B [11].
-In particular, the works by Gardner et al.
-[11] are note-worthy for curating the T4 dataset, containing roughly 3 M tables extracted and processed from the TabLib collection [8] and for its excellent results in the very low data-regime.
+TabLLM [tabllm], LIFT [lift], or TabuLa-8B [tabula8b].
+In particular, the works by tabula8b are note-worthy for curating the T4 dataset, containing roughly 3 M tables extracted and processed from the TabLib collection [tablib] and for its excellent results in the very low data-regime.
 3 Method
@@
 Any off-the-shelf embedding model can be used for this purpose – e.g.
-many are available based on the BERT architecture [5].
+many are available based on the BERT architecture [devlin-etal-2019-bert].
 As many values have to be embedded for each table, there is a natural trade-off between accuracy and speed.
-We settle for a comparably small and fast model such as all-MiniLM-L6-v2 [19, 34], since the amount of semantic meaning it captures is already much richer than conventional categorical encodings, and we defer to Section 5.1 for an experimental ablation.
+We settle for a comparably small and fast model such as all-MiniLM-L6-v2 [all-MiniLM-L6-v2, wang2020minilm], since the amount of semantic meaning it captures is already much richer than conventional categorical encodings, and we defer to Section LABEL:sec:ablation for an experimental ablation.
 Since the embeddings are potentially of a different dimension than the target embedding dimension $d$, we apply a learnable linear layer to the text embeddings.
@@
 recurring holidays), we embed each of the numbers representing day, month, and year separately and sum the three resulting vectors.
-In particular, this is more token efficient then encoding dates into multiple features as is common in established preprocessors such as the one used by AutoGluon.
+In particular, this is more token efficient than encoding dates into multiple features as is common in established preprocessors such as the one used by AutoGluon.
 Numerical:  As numbers do not contain semantic meaning beyond their value, we apply a one-dimensional encoding.
@@
 Second, we scale them to have zero mean and unit variance as is common.
-By Chebyshev’s inequality, this bounds resulting values to the interval $(-7.1,7.1)$, thus avoiding exploding gradients during training.
+By Chebyshev’s inequality, this bounds resulting values to the interval (7.1,7.1)(-7.1,7.1), thus avoiding exploding gradients during training.
 Finally, the resulting number is multiplied by a learnable vector and a bias is added.
@@
 The result is passed through a separate learnable linear layer to map to the correct target dimension and summed with the cell embedding.
-Note that all the above embeddings are fully equivariant under permutations of either rows or columns – a property that is often desirable in the tabular domain.
+After summation, the embeddings are normalized via layer normalization.
+Note that all the above embeddings are fully equivariant under permutations of either rows or columns.
 This equivariance makes the predictions more reproducible and robust and eliminates some of the need (such as the mapping of category to ID, or of column to ID) for bagging in models such as TabPFN.
-Semantic embeddings of cell values have been previously investigated in other works [35, 36, 22, 31], however, details of how the information is consumed differ in each implementation.
+Semantic embeddings of cell values have been previously investigated in other works [yak2023ingestables, ye2024crosstable, CARTE, PORTAL], however, details of how the information is consumed differ in each implementation.
 3.2 Backbone
@@
 Following TabPFN, cross-column attention has no masking, while cross-row attention is masked so that each row can only attend to the provided context.
-To increase the modularity of our code, the feedforward MLP block of the transformer encoder is repeated after each self-attention, so that “horizontal” and “vertical” blocks have exactly the same weights.
+To increase the modularity of our code, the feedforward MLP block of the transformer encoder is repeated after each self-attention block so that “horizontal” and “vertical” blocks have the same structure.
 For sizing, we use the nomenclature popularized by BERT, e.g.
-by calling “base” the variant with 12 layers and a hidden dimension of 768.
+calling “base” the variant with 12 layers and a hidden dimension of 768.
 However, the real number of (non-embedding) weights is twice that of a BERT model, due to the presence of both “horizontal” and “vertical” layers.
-Model weights can also be optionally shared between each instance of the transformer block, which consists of two interleaved attention layers.
-Such an architecture can be interpreted as a recurrent neural network (RNN), but unrolled in depth rather than in time.
+Model weights can be optionally shared between each instance of the transformer block, consisting of two interleaved attention layers.
+Such an architecture can be interpreted as a recurrent neural network, unrolled in depth rather than in time.
 Such an iterative structure results in parameter efficiency and the possibility of stacking more blocks in comparison to the traditional approach.
@@
 As further benefits, full equivariance is preserved, without the need to map classes to IDs, and the limitation on the number of classes supported for prediction is lifted, improving upon the constraint currently present in many tabular ICL approaches such as TabPFN, TabDPT, or TabICL.
-For this, we take inspiration from Polewczyk and Spinaci [26]: For each row, the final output corresponding to the target column is mapped by a two-layer MLP to a vector $x$.
+For this, we take inspiration from clustertabnet: For each row, the final output corresponding to the target column is mapped by a two-layer MLP to a vector $x$.
 We then build a matrix of shape $(n_{\mathrm{query}},n_{\mathrm{context}})$ by computing the cosine similarities between these vectors $x_{i}$, $x_{j}$ for each query row $i$ and context row $j$.
@@
 Induced Set Attention Blocks:  One of the main limitations of the standard multi-head attention is its quadratic complexity.
-While memory can be effectively compressed to linear [29, 4], that is not the case for runtime.
+While memory can be effectively compressed to linear [efficientattention, flashattention], that is not the case for runtime.
 This poses a significant challenge when dealing with large tables.
-Taking inspiration from [23, 20, 28], we experiment with methods to handle larger context more efficiently.
-Our proposed architecture replaces some of the multihead attention blocks within TabPFN with ISAB blocks from [23].
-Following the notation from [23], let $\mathrm{MAB}(X,Y)$ denote the original multihead attention block from the transformer architecture, followed by a feedforward block including skip connections and layer normalization.
-Let $I$ be so-called inducing points [23], an additional set of parameters of shape $(n,d)$ for an arbitrary $n$, with $d$ denoting the hidden dimension, and define
+Taking inspiration from [isab, perceiver, tabicl], we experiment with methods to handle larger context more efficiently.
+Our proposed architecture replaces some of the multihead attention blocks within TabPFN with ISAB blocks from [isab].
+Adopting the notation from [isab], let $\mathrm{MAB}(X,Y)$ denote the standard multihead attention block from the transformer architecture, where $X$ represents the queries and $Y$ serves as both the keys and values.
+This is followed by a feedforward block, a skip connection, layer normalization, and ultimately an intermediate block.
+Let $I$ be so-called inducing points [isab], an additional set of parameters of shape $(n,d)$ for an arbitrary $n$, with $d$ denoting the hidden dimension, and define
 $\mathrm{ISAB}(C\mathbin{+\mkern-10.0mu+}Q)=\mathrm{MAB}\big(C\mathbin{+\mkern-10.0mu+}Q,\mathrm{MAB}(I\mathbin{+\mkern-10.0mu+}Q,C\mathbin{+\mkern-10.0mu+}Q)\big),$ (2)
-[…]
+where $C$ and $Q$ denote context and query rows, respectively, and $\mathbin{+\mkern-10.0mu+}$ denotes concatenation.
+In each attention computation, we apply masking so that only context and inducing vectors are attended to.
+For a detailed overview of the architecture, refer to Appendix Figure LABEL:fig:ISAB_diagram.
 Since tables can have a very large number of rows but rarely exceed a few hundred columns, we apply the ISAB block only to cross-row attention.
@@
 4.1 Training and inference
-For pretraining, we use the T4 dataset [11].
+For pretraining, we use the T4 dataset [tabula8b].
 We discard tables with fewer than 150 rows, which leaves $2.18$ M tables with a median of 750 rows and 9 columns.
 We randomly select 1000 rows, then between 50 and 900 rows as query, and use the rest as context.
-Then we randomly select one target column, excluding all date columns, numerical columns with more than 50% NaN values, and other columns having more than 20% of unique values.
+Subsequently, we randomly select one target column, excluding all date columns, numerical columns with more than 50% NaN values, and other columns having more than 20% of unique values.
 Finally, we up-sample non-numeric columns to have roughly the same proportion of regression and classification tasks.
@@
 Hence, full training takes between 4 and 12 days depending on the number of steps.
-Using the default parameters $n=12$, $d=768$, $d_{ff}=3072$ would result in a total of 172M parameters, but due to weight sharing there are only 16M trainable parameters.
-We also experimented with curriculum learning, by further adding in a second step using the same training data as Ma et al.
-[24].
+Using the default parameters $n=12$, $d=768$, $d_{\textrm{ff}}=3072$ results in a total of 172 M parameters which is reduced to 16 M trainable parameters when using weight sharing.
+We also experimented with curriculum learning, by further adding in a second step using the same training data as tabdpt.
 This data has fewer tables, 123, but many more cells in each table with a median of 11 k rows and 34 columns.
 In this second step, we increased the number of rows used for training to 4000.
-We refer to Section 5.1 for an analysis of the impact of training data size on model performance, which we found to be crucial.
-At inference time, we apply 8-fold bagging, similar to the ensembling used by TabPFN.
-That is, from the original train split of a given evaluation dataset, we sample 8 times $c$ context rows with replacement, make a separate prediction with each collection of $c$ rows using the same model, and average the predictions (regression values or classification probabilities).
+We refer to Section LABEL:sec:ablation for an analysis of the impact of training data size on model performance, which we found to be crucial.
+At inference, we apply 8-fold bagging, similar to the ensembling used by TabPFN.
+That is, from the original train split of a given evaluation dataset, we sample 8 times $c$ context rows with replacement, make a separate prediction with each collection using the same model, and average the predictions (regression values or classification probabilities).
 While, during the default training, $c$ never surpasses 950, we can scale this during evaluation as much as memory permits.
 The combination with bagging allows ConTextTab to use up to $8c$ points as context.
-To limit runtime, in each bag, we also sample up to 500 columns if the original dataset has more than that.
+By default, we use a context size of $c=8192$ at inference.
+To limit runtime, in each bag, we sample up to 500 columns if the dataset has more.
 4.2 Evaluation
 Datasets:  We use a variety of tabular prediction datasets to evaluate and compare our approach to established baselines and other SOTA methods.
-Namely, we evaluate all models on the following benchmarks: OpenML-CC18 [1], a pure classification benchmark; OpenML-CTR23 [10], a pure regression benchmark; TALENT [37], a recently introduced diverse benchmark containing over 300 classification and regression benchmarks.
-Here, we focus on a subset containing 45 datasets that are representative of the overall performance of the baselines investigated in the original works, which we refer to as the TALENT-Tiny benchmark; TabReD [30], a small but challenging benchmark of large datasets representative of practical prediction tasks; and finally CARTE [22], a mixed classification and regression benchmark containing highly semantic features and few numerical ones.
+Namely, we evaluate all models on the following benchmarks: OpenML-CC18 [openmlcc18], a pure classification benchmark; OpenML-CTR23 [openmlctr23], a pure regression benchmark; TALENT [talent], a recently introduced diverse benchmark containing over 300 classification and regression benchmarks.
+Here, we focus on a subset containing 45 datasets that are representative of the overall performance of the baselines investigated in the original works, which we refer to as the TALENT-Tiny benchmark; TabReD [tabred], a small but challenging benchmark of large datasets representative of practical prediction tasks; and finally CARTE [CARTE], a mixed classification and regression benchmark containing highly semantic features and few numerical ones.
 Across all benchmarks, we evaluate 91 regression and 112 classification tasks, ranging from 5 to 3 k columns and 400 to roughly 400 k training examples.
@@
 For models that do not explicitly use a validation split, we concatenate the train and validation splits.
-We refer to Appendix B.2 for a visualization of the dataset statistics and further details.
+We refer to Appendix LABEL:app:dataset for a visualization of the dataset statistics and further details.
 Baselines:  We compare our approach to several established classical methods as well as recent ICL and other deep learning models.
-In particular, we compare with TabPFN [17], TabICL [28], and TabDPT [24] as the most recent table-native ICL methods.
-For all pretrained models, we use the latest available release and checkpoints as of May 2025.
-As the SOTA on the CARTE benchmark, we also evaluate the CARTE model proposed by [22].
-As another well-performing recent deep learning approach, we evaluate RealMLP [18].
-Note that both CARTE and RealMLP need to be finetuned or fitted for each dataset separately.
-We perform per-dataset hyperparameter tuning via random search for RealMLP as provided in the author’s implementation via the pytabkit package.
-Also using the implementation wrapper from pytabkit, we evaluate SOTA boosted tree baselines, namely XGBoost [2], LightBGM [21], and CatBoost [27] – both in the meta-tuned default variants (TD) [18] as well as via a Parzen-tree estimator-based hyperparameter optimization (HPO).
-Furthermore, we evaluate several common, non-tuned baselines from scikit-learn [25], namely naive, linear, and KNN estimators, as well as a random forest and histogram-based gradient boosting tree estimators.
-Finally, as the gold standard in tabular prediction, we evaluate AutoGluon [9], an AutoML solution stacking and ensembling many of the baselines outlined above.
-Please refer to Appendix B.1 for full details on the used baseline versions and parameters as well as details about the dataset preprocessing and encoding.
+In particular, we compare with TabPFN [tabpfnv2], TabICL [tabicl], and TabDPT [tabdpt] as the most recent table-native ICL methods.
+For all pretrained models, we use the latest available release and checkpoints as of July 2025.
+As the SOTA on the CARTE benchmark, we also evaluate the CARTE [CARTE].
+As additional well-performing recent deep learning approaches, we evaluate RealMLP [realmlp] and TabM [tabm].
+Furthermore, we evaluate SOTA boosted tree baselines, namely XGBoost [xgboost], LightBGM [lightgbm], and CatBoost [catboost].
+The tree baselines as well as RealMLP and TabM are evaluated in the default (D) or meta-tuned default variants (TD) [realmlp], as well as via Parzen-tree estimator-based hyperparameter optimization, either with a fixed holdout set (HPO) or ensembled via 5-fold inner cross-validation with per-split HPO (HPO, CV).
+For this, we use the implementation provided by pytabkit [realmlp].
+Additionally, we evaluate several common, non-tuned baselines from scikit-learn [sklearn], namely naive, linear, and KNN estimators, as well as a random forest and histogram-based gradient boosting tree estimators.
+Finally, as the gold standard in tabular prediction, we evaluate AutoGluon [autogluon], an AutoML solution stacking and ensembling many of the baselines outlined above.
+We refer to Appendix LABEL:app:baselines for full details on the used baseline versions and parameters as well as details about data preprocessing and encoding.
 Metrics:  Throughout, we show (mean) accuracy for classification tasks as well as (mean) $R^{2}$ score for regression tasks.
+Since we have observed some severe outliers in some regression tasks that greatly affected average scores, we perform a soft-clipping.
+That is, all negative $R^{2}$ scores are mapped to [1,0)[-1,0) via $\tanh$ while keeping positive scores untouched.
+This retains the relative ordering of scores while being smoothly differentiable and less distorting to the mean.
 As averaging across a large number of datasets with varying performance can blur relative performance across models, we also evaluate (mean) rank.
@@
 To reduce the impact of small noise, we count model performances as ties when their evaluation scores lie within 0.005 of each other.
+Furthermore, we report critical difference (CD) diagrams, win ratio matrices, and p-values enabling 1 vs 1 model comparisons in the Appendix.
 5 Results
-Table 1: Performance comparison across all evaluated benchmarks, depicting mean accuracy (Acc) for classification and $R^{2}$ score for regression tasks, in percent.
-Missing values, due to architectural limitations or failed evaluations, are denoted as N/A and excluded from the mean rank calculation.
-Model Name All CARTE OML-CC18 OML-CTR23 TabReD TALENT-Tiny Rank Rank Acc R2 Rank Acc Rank R2 Rank Acc R2 Rank Acc R2 AutoGluon 78.7 73.8 88.5 67.0 86.0 64.6 87.9 73.7 ConTextTab 2.20 1.49 76.9 72.2 2.79 86.8 3.29 72.9 3.88 85.4 63.4 2.30 87.7 76.1 TabPFN 2.66 5.08 72.4 65.0 2.67 87.0 2.63 74.9 3.00 85.6 63.8 1.73 87.3 75.1 LightGBM [HPO] 2.85 4.06 72.8 66.1 3.26 86.9 3.31 61.9 1.62 85.9 64.4 3.19 86.4 72.4 XGBoost [HPO] 3.16 4.47 72.6 65.7 4.00 86.3 3.17 71.4 1.75 86.0 64.2 3.24 86.2 72.6 CatBoost [HPO] 3.24 4.02 75.6 66.1 3.57 86.7 4.40 -44.4 2.00 85.8 63.7 3.76 86.0 72.1 HistGradBoost 3.46 5.08 72.5 64.8 3.62 86.1 4.80 51.2 1.75 85.9 63.9 3.62 86.3 67.6 RealMLP [TD] 3.91 8.39 70.2 59.6 3.18 87.2 3.80 45.7 1.25 86.0 64.6 4.08 86.2 71.5 Random forest 4.56 7.08 71.5 63.3 4.69 85.7 5.17 53.6 5.38 85.4 60.7 4.54 85.8 70.6 Naive 8.61 11.22 53.0 -1.8 9.39 47.0 9.69 -8.5 8.62 80.8 -0.6 9.97 53.4 -22.3 CARTE N/A 3.16 76.1 68.5 N/A N/A N/A N/A N/A N/A N/A N/A N/A N/A TabDPT N/A 5.02 72.7 65.1 2.71 87.8 2.46 72.4 N/A N/A 60.9 2.65 86.7 74.8 TabICL N/A N/A 72.5 N/A N/A 88.0 N/A N/A N/A 85.1 N/A N/A 87.4 N/A
-The main results, evaluating ConTextTab and the baselines as outlined in the previous section, are shown in Table 1.
+The main results, evaluating ConTextTab and baselines, are shown in Table 5.
 Here, we exclude AutoGluon from the overall rank and best-model comparison as it is ensembling and stacking a multitude of model architectures, making it difficult to highlight architectural strengths and weaknesses.
-Overall, our model achieves the highest rank across all evaluated datasets.
-This performance gain is statistically significant, as depicted in the critical difference diagram in Figure 2 (left).
-Our model performs particularly well for the semantically rich CARTE benchmark, highlighting the importance of incorporating semantic understanding into tabular ICL.
-Note that TabPFN, not incorporating semantics, performs worse than both tuned and non-tuned boosted trees in this particular instance.
-On the other end, while the absolute performance of ConTextTab for large datasets, e.g.
-within the TabReD benchmark, is not bad, it falls behind tuned boosted trees here, which still achieve SOTA, on par with AutoGluon.
-To investigate the dependence on dataset size more closely, we show the performance on the CARTE benchmark across varying subsampled sizes, ranging from 128 rows to the full dataset, in Figure 3.
+Overall, our model performs competitive across all non-semantic benchmarks while setting a new standard on the semantically rich CARTE benchmark, where it ranks best.
+In fact, this is statistically significant over all models but CatBoost and AutoGluon, both of which are extensively tuned and ensembled, as shown in the critical difference diagram for CARTE in Figure LABEL:fig:critical_difference_and_pretraining_size (left).
+Notably, ConTextTab is significantly better than other tabular ICL approaches, which highlights the importance of incorporating semantic understanding.
+On the other hand, TabPFN, not incorporating semantics, performs worse than tuned as well as untuned trees on CARTE.
+For details, we refer to Appendix LABEL:app:additional-results.
+On the other end, the absolute performance of ConTextTab for non-semantic benchmarks, while good, falls behind tuned and ensembled boosted trees as well as RealMLP, which achieves SOTA in terms of ranking across all evaluated datasets (albeit at a much higher training time and cost).
+Nevertheless, these differences in ranking are mostly not statistically significant.
+To this end we report per-benchmark CD diagrams as well as win-ratio and p-values for detailed model comparison in Appendix LABEL:app:additional-results-extended.
+Namely, ConTextTab does not perform significantly worse than the best-ranking model on all non-semantic benchmarks excluding OpenML-CTR23 (where it is however not significantly different to tuned ensembled trees).
+Hence, ConTextTab sets a new standard on semantically rich datasets while performing on-par to existing approaches on non-semantic ones.
+Nevertheless, the performance of AutoGluon generally indicates further headroom to improve single-model performance and tabular ICL in particular.
+Finally, tuned trees perform particularly well for large datasets as represented by the TabReD benchmark, even outperforming AutoGluon in some instances.
+Motivated by this, we investigate the dependence on dataset size more closely and the performance on the CARTE benchmark across varying subsampled sizes, ranging from 128 rows to the full dataset, in Figure LABEL:fig:carte-sub.
 ConTextTab consistently outperforms other models across all sample sizes and even surpasses AutoGluon for up to 2048 training samples.
 This highlights the strong capabilities of tabular ICL but also the need for further research into scaling these architectures to effectively deal with much larger context sizes as well as training datasets.
-We investigate the dataset-size dependent rank performance across all 203 datasets in Appendix A.1, further highlighting these challenges.
-Additional results, including additional baselines and win ratios, are given in Appendix A.2.
-Figure 2: Left: critical difference diagram between ConTextTab and several baselines, across all 203 evaluated datasets.
-Right: Impact of pretraining dataset size on validation accuracy and $R^{2}$ scores.
-Figure 3: Average rank, accuracy, and regression results on the CARTE benchmark across various data subsets, ranging from 128 rows to the full size.
-5.1 Ablation studies
-In the following, we discuss several experiments ablating aspects of our model reported in Table 2.
-Table 2: Multiple ablations as relative differences with respect to one of two base configurations, using binning (top) or one-dimensional embedding (bottom) for regression.
-Improvements over the base model are highlighted in bold.
-Experiment All CARTE OML-CC18 OML-CTR23 TabReD TALENT-Tiny Rank Rank Acc R2 Rank Acc Rank R2 Rank Acc R2 Rank Acc R2 base model (binning) 2.28 1.76 76.0 71.4 2.57 86.8 2.66 70.9 1.00 85.4 62.9 2.32 87.2 73.5 large size 2.76 2.18 +0.2 -0.2 4.01 -5.3 1.77 -0.7 3.75 -25.4 -13.5 1.84 -0.1 -0.0 medium size 5.44 8.45 -5.6 -3.7 4.76 -1.2 3.49 -1.0 5.75 -9.3 -3.5 4.41 -2.6 -2.4 small size 7.68 9.98 -2.9 -6.8 6.57 -3.0 8.17 -4.5 8.12 -1.7 -3.8 6.14 -2.7 -6.7 mini size 8.10 10.69 -4.2 -8.4 7.15 -4.0 8.06 -4.2 5.50 -0.3 -3.5 7.00 -3.7 -6.9 multilingual-e5-small 3.22 4.16 -0.3 -1.5 3.35 -0.4 2.77 -0.1 2.38 -0.2 -1.3 2.27 +0.7 -0.7 gte-multilingual-base 3.22 3.59 +0.0 -1.2 3.18 -0.3 3.60 -1.4 1.62 -0.2 -0.9 2.76 +0.4 -1.3 curriculum learning 2.28 1.51 +1.1 +0.1 3.10 -0.2 1.80 -0.1 1.38 +0.4 -0.2 2.38 -1.0 +0.4 clustering N/A N/A -0.1 N/A 3.82 -1.5 N/A N/A N/A -0.6 N/A N/A -1.3 N/A w/o bagging 2.82 2.76 -0.4 -0.4 3.26 -0.7 3.03 -0.4 1.00 +0.1 -0.5 2.24 -0.6 -0.0 ISAB 5.48 6.16 +0.3 -3.1 5.57 -2.1 5.51 -0.9 3.88 -0.2 -2.7 4.70 -0.5 -6.1 non-shared weights 4.36 5.61 -0.1 -2.6 4.29 -0.9 3.09 +0.8 5.88 -1.0 -2.6 3.65 -0.7 -1.2 base model (1-dim) 2.00 1.80 76.7 72.0 2.36 86.9 2.06 65.0 1.75 85.3 63.3 1.54 87.9 75.8 0.1% clipping 2.03 1.80 0.0 0.0 2.36 0.0 2.37 -0.3 1.38 0.0 +0.2 1.54 0.0 0.0 0.5% clipping 2.08 1.76 0.0 0.0 2.32 -0.1 2.60 +4.0 1.38 +0.1 +0.1 1.70 +0.1 -0.2 2% clipping 2.97 2.88 0.0 -0.5 2.43 -0.1 5.03 +5.8 1.75 +0.1 -0.5 2.43 0.0 -2.0 binning 4.04 4.69 -0.7 -0.6 3.21 -0.1 6.34 +5.9 1.12 +0.1 -0.3 3.24 -0.7 -2.3 context=20000 2.06 1.39 +0.3 +0.6 2.62 -0.2 1.66 +4.2 3.25 -0.4 -0.5 2.00 -0.4 +0.5 context=10000 2.00 1.61 +0.3 +0.3 2.64 -0.2 1.57 +4.2 1.75 0.0 -0.1 1.76 -0.2 +0.5 context=4096 2.81 3.55 -0.1 -0.6 3.03 -0.3 2.11 +4.0 2.25 -0.3 -0.1 2.16 -0.4 +0.4 context=2048 4.11 6.41 -1.0 -1.9 3.61 -0.4 3.06 +3.6 2.50 -0.1 -0.7 3.24 -0.9 +0.1 context=1024 5.49 9.10 -1.3 -3.8 4.33 -0.5 4.60 +2.7 2.62 -0.4 -1.6 4.24 -1.3 -0.6 fine-tuning 2.60 2.51 +0.3 +0.1 3.1 +0.2 1.94 +7.2 2.62 +0.2 -0.3 2.38 -0.1 +0.8
-Model size:  As expected, decreasing from “base” to smaller models impacts performance.
-On the other hand, increasing to “large”, thus doubling the number of layers and increasing the hidden dimension, has no statistically significant impact on performance.
-As observed in Figure 2 (right), increasing the amount of training data leads to stagnating validation performance gains.
-We hypothesize that, in the current setup, our model is likely limited by the amount or diversity of data available with T4.
-Context size:  The trained models can scale the input context size at inference.
-As expected, longer context improves results in a monotonic way, in exchange for longer computation time.
-In practice, the effects are stronger on larger datasets, simply because shorter context models (even with bagging) only get to see a subset of the training data.
-We stress again that this is entirely extrapolating outside of the pretraining setup, where the largest observed context has size 950.
-Nevertheless, with larger datasets and context, the model performance saturates, requiring further research in how to increase the capacity of tabular ICL.
-Training data:  We ran training on randomly chosen subsets of the T4 dataset of $10^{n}$ tables, $n\in\{1,\dots,6\}$.
-For this experiment we report the validation scores (on the validation held-out subset of the CARTE dataset) in Figure 2 (right).
-The results suggest that at least 100 k of the tables are needed in order to train a model of state-of-the-art quality.
-While exact numbers may vary when using other datasets or model sizes, this gives a high-level estimate of the amount of required data, giving indication of why our model could not profit from a second step of curriculum learning on the much smaller dataset used by Ma et al.
-[24].
-Potentially, training is still limited by data amount or diversity, e.g.
-missing longer and wider tables, likely required to further scale tabular ICL.
-Classification target:  We compare the default cross-entropy loss for classification with the clustering approach described in Section 3.4, allowing to retain the semantics of the target labels, lifting the limitation on the number of target classes as present in TabPFN or TabICL.
-However, most evaluation datasets only contain non-descriptive labels, and the number of classes is limited.
-Hence, the potential of clustering cannot emerge, and the default maintains a modest edge.
-We believe that semantically richer high-cardinality benchmarks are required to spur research in this direction.
-ISAB block:  We observed comparable results on classification tasks and a modest drop in performance on regression tasks, while substantially enhancing efficiency by eliminating quadratic time complexity and reducing runtime by a factor of ten for large contexts.
-Bagging:  As expected, using bagging at inference time leads to a consistent improvement of results.
-At the same time, in absolute terms, the gap is rather marginal, probably thanks to the fact that ConTextTab gets rid of most sources of non-equivariance in the architecture (e.g.
-category indexing and positional encoding).
-As a result, depending on the use case, this allows to drop bagging, or reduce the number of bags, without compromising performance significantly.
-(No) weight sharing:  We observed a slight performance degradation, which may be due to either an insufficient training duration or suboptimal selection of training hyperparameters.
-Further ablations:  Other changes (the used sentence embedding model, longer context in training, target clipping for regression, as well as benchmark-dependent fine-tuning) did not result in significant performance gains.
-Therefore, we refer to the Appendix A.3 for further discussion of these ablations.
-6 Conclusions
-Limitations and future work:  While achieving SOTA results across the investigated datasets, we observe several limitations of our proposed approach.
-One drawback of using real-world data for training is the possibility of contamination, e.g.
-the presence of evaluation tasks in the training corpus.
-Since the CARTE benchmark is our focus, we conducted a contamination study, using the column name and cell value embeddings created by a text embedder and matching similarities.
-We did not find any contamination of CARTE in T4.
-For contamination of OpenML datasets, we refer to the original study by [11].
-Either way, as a single table is only seen a few times during training, we believe that memorization is likely not a practical problem when training on real-world data.
-Generally, all investigated table-native ICL models, e.g.
-ours, TabPFN, TabDPT, and TabICL, fail to scale their performance to very large datasets.
-Increasing the context length did not resolve these issues.
-Using local context, as done in TabDPT, might overcome such limitations but we were unable to evaluate TabDPT on TabReD as it could not handle large tables.
-In the large-data case, conventional methods, in particular when stacked and ensembled via AutoGluon, still perform best.
-Overcoming this remains one of the major research challenges for tabular foundation models.
-We observe little gain when training on larger data corpora, e.g.
-increasing the pretraining subset of T4 or adding tables from the AutoML benchmark.
-Together with the investigated scaling of our model, we believe that more diverse real-world data is needed to fully unlock semantic understanding at scale.
-While models trained on synthetic data do not face this issue in principle, how to incorporate semantic alignment in these models remains an open question.
-Similarly for evaluation, semantically rich benchmarks are scarce.
-While CARTE includes rich semantic features, all classification targets are binary, containing no semantic information.
-Furthermore, it contains mostly small to medium sized datasets, in particular in terms of the number of features.
-Hence, we argue that more diverse, semantically rich data is required – both at scale for pretraining larger models with longer context, as well as curated for evaluation.
-Finally, our investigation to utilize target semantics and lift the strict class limit at inference prevalent in current tabular ICL, by using a dynamic supervised clustering approach, did not result in performance gains.
-This might relate to the lack of semantically meaningful targets in current tabular benchmarks as well as the limited target cardinality across them.
-Summary:  We have introduced ConTextTab – a context aware table-native in-context learner trained on real-world data for both classification and regression tasks.
-Evaluated across a wide range of tabular benchmarks – including both classification and regression tasks – our model achieves state-of-the-art results significantly outperforming existing table-native ICL approaches such as TabPFN, in particular on the semantically rich CARTE benchmark and low-data regime.
-We outline future research efforts required to further improve semantic understanding in tabular foundation models, in particular the need for more diverse semantically rich pretraining as well as evaluation data.
-Acknowledgements
-We would like to thank Johannes Hoehne, Johannes Hoffart, and Markus Kohler for their insightful comments and suggestions throughout the development of this work, which have greatly contributed to shaping its direction and quality.
-Appendix A Further Results
-A.1 Relation between dataset size and model performance
-We plot the average rank of each model as a function of the dataset size (expressed in number of rows) across all evaluated datasets in Figure 4.
-Note that TabICL is missing from this comparison as it only handles classification tasks.
-We can observe that, as expected, ConTextTab and TabPFN excel in the low data regime, while AutoGluon takes the lead when enough data is available.
-In the very low data regime below 1000 training rows, TabPFN performs best, however our model surpasses its performance for larger datasets with more than 1000 rows.
-Overall, ConTextTab remains competitive with AutoGluon until roughly 10 000 rows, possibly as a result of the limitation in both the inference context size and the context size seen during training.
-After 10 000 rows, gradient boosting methods, as well as RealMLP, start surpassing the performance of ConTextTab as well as TabPFN.
-Overall, this indicates the need of further research for tabular ICL to handle larger amounts of training data, but also the availability of more diverse benchmarks, covering larger datasets, as the current evaluation is dominated by datasets with less than 10 k rows and less than 100 columns, see Figure 6.
-Figure 4: Relation between number of training dataset rows and performance, obtained as a LOWESS regression in the plane $\log(n_{\mathrm{rows}},\textnormal{rank})$.
-The confidence bands are the 80% confidence intervals obtained via bootstrapping.
-A.2 Extended results
-Including further baselines not depicted in the main paper, Table 3 shows all collected evaluations across the investigated benchmarks.
-Furthermore, we show the win ratio confusion matrix and averages of the investigated models in Figure 5.
-(a) Win ratio matrix with Model A wins over Model B.
-(b) Average Win ratio of Model A against all others.
-Figure 5: Win ratio confusion matrix and average of the investigated models across all 203 datasets.
-Wins are calculated based on accuracy on classification and $R^{2}$ on regression datasets.
-Ties are not counted as wins.
-Models are sorted by descending overall rank.
-Note that CARTE, TabDPT, RealMLP [HPO], and TabICL were not successfully evaluated on all datasets and only compared on evaluated ones, resulting in a smaller support and skewing their win rates.
-Therefore, they are separately grouped last, not used in averaging, and only shown for completeness.
-Table 3: Performance comparison across all evaluated benchmarks, depicting mean accuracy (Acc) for classification and $R^{2}$ score for regression tasks, in percent.
-Missing values, due to architectural limitations or failed evaluations, are denotes as N/A and excluded from the mean rank calculation.
-Model Name All CARTE OML-CC18 OML-CTR23 TabReD TALENT-Tiny Rank Rank Acc R2 Rank Acc Rank R2 Rank Acc R2 Rank Acc R2 Autogluon 2.48 1.55 78.7 73.8 2.61 88.5 4.06 67.0 1.75 86.0 64.6 2.16 87.9 73.7 ConTextTab 3.53 2.24 76.9 72.2 4.08 86.8 4.37 72.9 4.62 85.4 63.4 3.22 87.7 76.1 TabPFN 4.27 7.25 72.4 65.0 3.61 87.0 3.43 74.9 4.5 85.6 63.8 2.16 87.3 75.1 LightGBM [HPO] 4.59 5.41 72.8 66.1 4.56 86.9 4.63 61.9 1.75 85.9 64.4 4.11 86.4 72.4 CatBoost [HPO] 4.99 5.33 75.6 66.1 4.6 86.7 6.63 -44.4 2.12 85.8 63.7 4.32 86.0 72.1 CatBoost [TD] 4.99 6.35 75.1 65.6 4.96 86.4 4.51 67.2 1.25 85.9 64.3 4.43 85.9 74.8 LightGBM [TD] 5.05 7.02 73.1 64.9 4.01 86.8 6.29 39.7 2.25 85.7 63.2 3.81 86.3 72.6 XGBoost [HPO] 5.05 5.84 72.6 65.7 5.74 86.3 4.34 71.4 1.75 86.0 64.2 4.03 86.2 72.6 XGBoost [TD] 5.52 8.24 72.3 64.4 4.03 87.0 6.54 63.7 2.5 85.6 62.8 4.35 86.0 72.7 HistGradBoost 5.81 7.04 72.5 64.8 5.11 86.1 7.14 51.2 2.25 85.9 63.9 5.0 86.3 67.6 RealMLP [TD] 6.52 12.2 70.2 59.6 4.62 87.2 5.14 45.7 1.12 86.0 64.6 4.86 86.2 71.5 Random forest 7.59 9.82 71.5 63.3 6.57 85.7 7.77 53.6 7.25 85.4 60.7 6.38 85.8 70.6 Linear 13.46 16.55 62.7 19.1 11.17 80.9 14.2 43.8 12.12 80.8 -4269.6 13.27 80.5 41.3 KNN [k=5] 13.78 16.25 65.5 34.3 12.06 81.7 13.57 14.1 13.5 78.7 -15.4 14.0 80.3 60.0 Naive 16.35 18.04 53.0 -1.8 15.28 47.0 16.4 -8.5 12.62 80.8 -0.6 16.89 53.4 -22.3 CARTE N/A 4.65 76.1 68.5 N/A N/A N/A N/A N/A N/A N/A 7.38 84.4 71.1 TabDPT N/A 6.63 72.7 65.1 3.74 87.8 3.26 72.4 N/A N/A 60.9 3.73 86.7 74.8 RealMLP [HPO] N/A 11.78 70.4 60.9 4.08 87.3 4.51 53.2 N/A N/A N/A 4.0 86.3 72.1 TabICL N/A N/A 72.5 N/A N/A 88.0 N/A N/A N/A 85.1 N/A N/A 87.4 N/A
-A.3 Further ablations
-In the following, we discuss in more detail the ablations as presented in Table 2.
-Text embedding model:  We test the performance of two more recently released sentence embedding models, namely Multilingual e5 small [33] and gte-multilingual-base [38].
-However, we cannot observe significant impact on the final results over the default all-MiniLM-L6-v2 sentence embedder.
-This potentially indicates the necessity of larger, semantically rich training data or further research into aligning these sentence embedders with a tabular foundation model.
-Curriculum training with longer context:  In this setup, after training the model on T4 with a context size of up to 950, training is extended by including also the dataset used by Ma et al.
-[24], collected from the OpenML AutoML benchmark, and extending the context size up to 4000 rows.
-Data is sampled from either T4 with 80% probability or the dataset from by Ma et al.
-[24] with 20% probability.
-The resulting model shows some improvements on a few tasks, but by statistically insignificant amounts.
-Regression target and clipping:  When using a one-dimensional encoding for numerical values, we can choose to clip the tails of the distribution for better stability.
-This is needed in pretraining due to the noisy data.
-In practice, whether to do this on downstream tasks can be chosen depending on outlier analysis of the data itself.
-We test alternative levels of clipping: no clipping (default), clipping at 0.1%, 0.5% and 2% quantiles.
-We also compare this to using binning instead of one-dimensional embeddings.
-We find that both clipping by 2% and binning cause a slight reduction in quality (since we are using $n=16$ bins, note that binning also incurs in clipping at $\frac{1}{32}\approx 3\%$).
-The only exception is in the OpenML-CTR23 benchmark: it contains one dataset with extreme outliers where no model achieves a positive $R^{2}$ score, and adding clipping there limits degenerate scores.
-Impact of fine-tuning:  The training process can be easily adapted to fine-tuning on a specific collection of datasets or even a single task.We experimented with fine-tuning on (the training partition of) all the 203 datasets at the same time.However, we observe that, if during this fine-tuning we fix the target column to always be the one corresponding to the prediction task to be evaluated, we rapidly encountered overfitting and prediction scores degrade.
-This might be a similar problem to the scaling law of training data, as also observed by Ma et al.
-[24].
-In the present situation, we can avoid overfitting by selecting the target column randomly with the same selection procedure described in Section 4.1.
-However, as shown in Table 2, the gain is rather limited.
-Appendix B Experimental Setup – Details
-B.1 Baselines
-Data preprocessor:  Evaluating models across a multitude of datasets can be tricky.
-Datasets may have inconsistent data type annotations, such as categories represented as strings or categorical data types, covering low- and high-cardinality categorical features, date, time, or datetime instances, free text, boolean values and more.
-Most models, however, require numerical input to process and handle non-numerical data types or missing values differently.
-To unify our evaluation, we implemented a configurable default feature encoder built on the AutoMLPipelineFeatureGenerator from AutoGluon [9] which we found to be very versatile and robust.
-In particular, the encoder natively handles low- and high-cardinality categorical data, free text (to some extent), as well as datetime encoding.
-For flexibility and compatibility with a multitude of models, we extended the default implementation to cover the following options that can be adapted to the capabilities of the baseline model at hand:
-  • Whether to convert booleans to string/categorical values
-  • Whether to not encode string/categorical values for models that natively handle them, such as CatBoost
-  • Whether to scale numerical data via quantile scaling with a normal distribution as target
-  • Whether to drop constant features
-  • Whether to impute missing values, extending the standard imputation (using most frequent categories and mean for numerical data) to bools and datetime data types
-As deault, we choose to convert booleans, encode categoricals via ordinal encoding, scale numerical data, drop constant columns and impute missing values.
-Below, we describe for which baselines the default values are changed or when other types of feature encodings are used.
-TabPFN:  We use the model from the official Python tabpfn package with version 2.0.8 together with the tabpfn-extensions package version 0.1.0 at commit d44606e35f89e18b6bc4c4a2eef2f46918c4302e of the Git repository [^2] https://github.com/PriorLabs/tabpfn-extensions.git as the PyPi release is not up-to-date.
-Naturally, we use TabPFNClassifier and TabPFNRegressor for classification and regression tasks, respectively, using default parameters for both.
-In particular, TabPFNClassifier uses an ensemble of 4 and TabPFNRegressor an ensemble of 8 estimators.
-We combine the classification estimator with the ManyClassClassifier extension with a redundancy factor of 4 to enable classification beyond the native 10-class limit of TabPFN which is required for the evaluation of some of the 203 evaluated datasets.
-For datasets larger than the native 10 k limit of TabPFN, we sample a random 10 k subset of the training split.
-This affects 66 out of the 203 evaluated datasets.
-For datasets with more than the 500 feature limit that TabPFN was trained with, we select a random subsample of 500 features.
-This affects 12 out of 203 evaluated datasets.
-While this is not optimal, and post-hoc ensembling as well as a random forest preprocessing is recommended by the authors [17], these extensions cannot be combined with the many-class extension required to predict beyond the 10-class limit of the native TabPFN model.
-Hence, we cannot evaluate TabPFN with the post-hoc ensembling or random forest extension.
-As we found the native feature encoder of TabPFN to not work across all evaluated datasets, we use our standard feature encoder (see above), encoding categorical columns, scaling numerical values, dropping constant columns, and imputing missing values.
-As this procedure should be very similar to the TabPFN-native encoder, we anticipate this deviation to affect the results only insignificantly if at all.
-TabICL:  We use the latest model weights tabicl-classifier-v1.1-0506.ckpt from the recent 0.1.2 version of the official tabicl package.
-This updated variant is an improved checkpoint over the one reported in the original works [28].
-As TabICL solely supports classification tasks, we exclude it from the overall mean rank evaluation.
-For encoding, we use our default encoder, but do not scale numericals, do not drop constant values, and do not impute missing ones as it is natively handled by the model.
-TabDPT:  We use the model from the official GitHub repository [^3] https://github.com/layer6ai-labs/TabDPT.git at the most recent 1.1.0 release and tabdpt1_1.pth model checkpoint.
-Naturally, we use TabDPTClassifier and TabDPTRegressor for classification and regression tasks, respectively, using default parameters for both.
-Throughout, we evaluate the model with a (local) context size of 2048 which is the best-performing one in the original works [24].
-However, evaluation failed for some datasets due to an error in the original code leading to empty predictions for very large datasets in the TabReD benchmark.
-Here, we use our default encoder, scaling numericals, dropping constant values, and imputing missing ones.
-CARTE:  We use the model provided in the official Python carte-ai package with version 0.0.26.
-We use CARTEClassifier and CARTERegressor with default parameters for classification and regression tasks, respectively We treat binary classification tasks as 2-class multi-class classification and hence set loss=‘‘categorical_crossentropy’’ for the classifacation estimator.
-With CARTE, we use our default preprocessor to convert bool values and datetime instances and to impute missing values, but otherwise rely on the Table2GraphTransformer provided in the reference implementation.
-Pytabkit models:  We use the pytabkit implementation wrapper for evaluating RealMLP, XGBoost, LightGBM, and CatBoost.
-We evaluate all models both in the tuned-defaults (TD) variant proposed by Holzmüller et al.
-[18] as well as hyperparameter-optimized (HPO).
-However, the HPO version did run into OOM issues for RealMLP, even when running on a H100 with 96 GB of VRAM which is why we display the TD variant by default.
-In particular, for RealMLP (TD), we use RealMLP_TD_Classifier and RealMLP_TD_Regressor for classification and regression tasks, respectively.
-For RealMLP (HPO), we use RealMLP_HPO_Classifier and RealMLP_HPO_Regressor for classification and regression tasks, respectively, conducting the default 50 rounds of random search HPO.
-For XGBoost (TD), LightGBM (TD), and CatBoost (TD), we use XGB_TD_Classifier, XGB_TD_Regressor, LGBM_TD_Classifier, LGBM_TD_Regressor, CatBoost_TD_Classifier, and CatBoost_TD_Regressor for classification and regression tasks, respectively.
-For the HPO-variants, we use the HPO_TPE versions of the estimators, performing Parzen-tree based HPO with 50 rounds using the search space as defined by [14].
-LightGBM and XGBoost are evaluated on 8-core CPU machines with 64 GB of RAM, whereas CatBoost and RealMLP are evaluated on H100 GPUs with 96 GB of VRAM.
-Note that CatBoost evaluation on CPU was too slow to evaluate at scale, in particular in the HPO variant.
-However, there are known issues with the GPU implementation of CatBoost which might degrade performance [^4] See https://github.com/catboost/catboost/issues/1408..
-We were not able to observe systematically worse results on those datasets on which we were also able to evaluate the CPU variant.
-Hence, for consistency, we present results for the GPU variant throughout.
-Throughout, we use our default encoder, scaling numericals, dropping constant values, and imputing missing ones. For all models but CatBoost, we perform ordinal encoding of categoricals.
-Sklearn models:  We use several standard baseline models from scikit-learn [25], combining them with the default preprocessor as outlined above.
-Across all scikit-learn baselines, preprocessing only differs in missing value imputation, depending on the model’s capability to handle missing values natively.
-Throughout, evaluation is performed using scikit-learn v1.5.2.
-For the naive predictor, we use the DummyClassifier and DummyRegressor to predict the most frequent, respectively mean value of the train splits as the naive majority baseline.
-For the linear predictor, we use the LogisticRegression and LinearRegression for classification and regression tasks, respectively, using default hyperparameters.
-For the KNN predictor, we use the KNeighborsClassifier and KNeighborsRegressor for classification and regression tasks, respectively, using default hyperparameters and $k=5$ nearest neighbors.
-For the random forest predictor, we use the RandomForestClassifier and RandomForestRegressor for classification and regression tasks, respectively, using default hyperparameters.
-The model handles missing values natively.
-Finally, for the histogram-based gradient boosted tree predictor, we use the HistGradientBoostingClassifier and HistGradientBoostingRegressor for classification and regression tasks, respectively, using default hyperparameters.
-The model handles missing values natively.
-AutoGluon:  Throughout, we use AutoGluon v1.2 and its TabularPredictor without custom preprocessing.
-We use the best_quality preset and set a per-dataset time limit of 4 h.
-Otherwise, parameters are left at their default values.
-For all datasets, evaluation is executed on a single 16-core machine with 128 GB of RAM and no GPU.
-B.2 Datasets
-Full details of all used dataset and benchmarks are provided in Table LABEL:tab:all-dataset-tables. The row and column count statistics are further visualized in Figure 6.
-We extracted all datasets from their original source and performed a custom stratified train-validation-test split with a 70-10-20 ratio.
-For classification tasks, the target column is used for stratification.
-For regression tasks, we perform stratification on the binned target column, binning it into 5 quantiles using the qcut method from the pandas library.
-Otherwise, we do not perform any alterations on the data.
-Models not using a specific validation procedure are provided with the concatenated train and validation split for training.
-Figure 6: Column and row distribution of the evaluated benchmark datasets.
-Table 4: Details of all used benchmark datasets.
-Benchmark Table Num. rows Num. cols Target type
-carte anime_planet 11512 11 regression
-carte babies_r_us 4068 5 regression
-carte beer_ratings 2557 20 regression
-carte bikedekho 3828 8 regression
-carte bikewale 7193 8 regression
-carte buy_buy_baby 8574 5 regression
-carte cardekho 30250 17 regression
-carte chocolate_bar_ratings 2070 9 classification
-carte clear_corpus 3779 27 regression
-carte coffee_ratings 1661 9 classification
-carte company_employees 8752 8 regression
-carte employee_remuneration 28316 5 regression
-carte employee_salaries 7368 9 regression
-carte fifa22_players 14468 19 regression
-carte filmtv_movies 32964 10 regression
-carte journal_jcr 7692 10 regression
-carte journal_sjr 22344 10 regression
-carte jp_anime 12428 17 regression
-carte k_drama 991 13 regression
-carte michelin 5465 8 classification
-carte mlds_salaries 8364 9 regression
-carte movies 5779 14 regression
-carte museums 9173 17 regression
-carte mydramalist 2720 14 regression
-carte nba_draft 1335 7 classification
-carte prescription_drugs 1371 8 regression
-carte ramen_ratings 3267 5 classification
-carte roger_ebert 2149 8 classification
-carte rotten_tomatoes 5726 14 regression
-carte spotify 32879 19 classification
-carte us_accidents_counts 18098 7 regression
-carte us_accidents_severity 17324 11 classification
-carte us_presidential 15885 7 regression
-carte used_cars_24 4734 9 regression
-carte used_cars_benz_italy 13112 8 regression
-carte used_cars_dot_com 3207 11 regression
-carte used_cars_pakistan 58124 8 regression
-carte used_cars_saudi_arabia 4405 11 regression
-carte videogame_sales 13128 6 regression
-carte whisky 1449 7 classification
-carte wikiliq_beer 10768 10 regression
-carte wikiliq_spirit 9820 8 regression
-carte wina_pl 1797 16 regression
-carte wine_dot_com_prices 12203 10 regression
-carte wine_dot_com_ratings 3276 10 regression
-carte wine_enthusiasts_prices 96780 10 regression
-carte wine_enthusiasts_ratings 103976 10 regression
-carte wine_vivino_price 11067 8 regression
-carte wine_vivino_rating 11067 9 regression
-carte yelp 51692 13 classification
-carte zomato 49152 9 classification
-openml_cc18 adult 39073 15 classification
-openml_cc18 analcatdata_authorship 672 71 classification
-openml_cc18 analcatdata_dmft 637 5 classification
-openml_cc18 balance_scale 500 5 classification
-openml_cc18 bank_marketing 36168 17 classification
-openml_cc18 banknote_authentication 1097 5 classification
-openml_cc18 bioresponse 3000 1777 classification
-openml_cc18 blood_transfusion_service_center 598 5 classification
-openml_cc18 breast_w 559 10 classification
-openml_cc18 car 1382 7 classification
-openml_cc18 churn 4000 21 classification
-openml_cc18 cifar_10 48000 3073 classification
-openml_cc18 climate_model_simulation_crashes 432 19 classification
-openml_cc18 cmc 1178 10 classification
-openml_cc18 cnae_9 864 857 classification
-openml_cc18 connect_4 54045 43 classification
-openml_cc18 credit_approval 552 16 classification
-openml_cc18 credit_g 800 21 classification
-openml_cc18 cylinder_bands 432 38 classification
-openml_cc18 devnagari_script 73600 1025 classification
-openml_cc18 diabetes 614 9 classification
-openml_cc18 dna 2548 181 classification
-openml_cc18 dresses_sales 400 13 classification
-openml_cc18 electricity 36249 9 classification
-openml_cc18 eucalyptus 588 20 classification
-openml_cc18 fashion_mnist 56000 785 classification
-openml_cc18 first_order_theorem_proving 4894 52 classification
-openml_cc18 gesturephasesegmentationprocessed 7898 33 classification
-openml_cc18 har 8239 562 classification
-openml_cc18 ilpd 466 11 classification
-openml_cc18 internet_advertisements 2623 1559 classification
-openml_cc18 isolet 6237 618 classification
-openml_cc18 jm1 8708 22 classification
-openml_cc18 jungle_chess_2pcs_raw_endgame_complete 35855 7 classification
-openml_cc18 kc1 1687 22 classification
-openml_cc18 kc2 417 22 classification
-openml_cc18 kr_vs_kp 2556 37 classification
-openml_cc18 letter 16000 17 classification
-openml_cc18 madelon 2080 501 classification
-openml_cc18 mfeat_factors 1600 217 classification
-openml_cc18 mfeat_fourier 1600 77 classification
-openml_cc18 mfeat_karhunen 1600 65 classification
-openml_cc18 mfeat_morphological 1600 7 classification
-openml_cc18 mfeat_pixel 1600 241 classification
-openml_cc18 mfeat_zernike 1600 48 classification
-openml_cc18 miceprotein 864 78 classification
-openml_cc18 mnist_784 56000 785 classification
-openml_cc18 nomao 27572 119 classification
-openml_cc18 numerai28_6 77056 22 classification
-openml_cc18 optdigits 4496 65 classification
-openml_cc18 ozone_level_8hr 2027 73 classification
-openml_cc18 pc1 887 22 classification
-openml_cc18 pc3 1250 38 classification
-openml_cc18 pc4 1166 38 classification
-openml_cc18 pendigits 8793 17 classification
-openml_cc18 phishingwebsites 8844 31 classification
-openml_cc18 phoneme 4323 6 classification
-openml_cc18 qsar_biodeg 844 42 classification
-openml_cc18 satimage 5144 37 classification
-openml_cc18 segment 1848 17 classification
-openml_cc18 semeion 1274 257 classification
-openml_cc18 sick 3017 30 classification
-openml_cc18 spambase 3680 58 classification
-openml_cc18 splice 2552 61 classification
-openml_cc18 steel_plates_fault 1552 28 classification
-openml_cc18 texture 4400 41 classification
-openml_cc18 tic_tac_toe 766 10 classification
-openml_cc18 vehicle 676 19 classification
-openml_cc18 vowel 792 13 classification
-openml_cc18 wall_robot_navigation 4364 25 classification
-openml_cc18 wdbc 455 31 classification
-openml_cc18 wilt 3871 6 classification
-openml_ctr23 abalone 3341 9 regression
-openml_ctr23 airfoil_self_noise 1202 6 regression
-openml_ctr23 auction_verification 1634 8 regression
-openml_ctr23 brazilian_houses 8553 10 regression
-openml_ctr23 california_housing 16512 9 regression
-openml_ctr23 cars 643 18 regression
-openml_ctr23 concrete_compressive_strength 824 9 regression
-openml_ctr23 cps88wages 22524 7 regression
-openml_ctr23 cpu_activity 6553 22 regression
-openml_ctr23 diamonds 43152 10 regression
-openml_ctr23 energy_efficiency 614 9 regression
-openml_ctr23 fifa 15342 29 regression
-openml_ctr23 forest_fires 413 13 regression
-openml_ctr23 fps_benchmark 19699 44 regression
-openml_ctr23 geographical_origin_of_music 847 117 regression
-openml_ctr23 grid_stability 8000 13 regression
-openml_ctr23 health_insurance 17817 12 regression
-openml_ctr23 kin8nm 6553 9 regression
-openml_ctr23 kings_county 17290 22 regression
-openml_ctr23 miami_housing 11145 16 regression
-openml_ctr23 moneyball 985 15 regression
-openml_ctr23 naval_propulsion_plant 9547 15 regression
-openml_ctr23 physiochemical_protein 36584 10 regression
-openml_ctr23 pumadyn32nh 6553 33 regression
-openml_ctr23 qsar_fish_toxicity 726 7 regression
-openml_ctr23 red_wine 1279 12 regression
-openml_ctr23 sarcos 39146 22 regression
-openml_ctr23 socmob 924 6 regression
-openml_ctr23 solar_flare 852 11 regression
-openml_ctr23 space_ga 2485 7 regression
-openml_ctr23 student_performance_por 519 31 regression
-openml_ctr23 superconductivity 17010 82 regression
-openml_ctr23 video_transcoding 55027 19 regression
-openml_ctr23 wave_energy 57600 49 regression
-openml_ctr23 white_wine 3918 12 regression
-tabred acquire_valued_shoppers_challenge 133602 117 classification
-tabred cooking_time 278338 196 regression
-tabred delivery_eta 313589 227 regression
-tabred home_credit_credit_risk_model_stability 325663 701 classification
-tabred homesite_quote_conversion 244458 302 classification
-tabred maps_routing 219994 991 regression
-tabred sberbank_russian_housing_market 23674 396 regression
-tabred tabred_weather 382955 111 regression
-talent_tiny ailerons 11000 41 regression
-talent_tiny breast_w 31492 10 classification
-talent_tiny cmc 44236 10 classification
-talent_tiny dis 3017 30 classification
-talent_tiny eye_movements_bin 6086 21 classification
-talent_tiny fried 32614 11 regression
-talent_tiny house_16h_reg 18227 17 regression
-talent_tiny ibm_employee_performance 1176 31 classification
-talent_tiny jungle_chess_2pcs_raw_endgame_complete 35855 7 classification
-talent_tiny kaggle_bike_sharing_demand_challange 8708 10 regression
-talent_tiny kc1 1687 22 classification
-talent_tiny kin8nm 6553 9 regression
-talent_tiny law_school_admission_bianry 16640 12 regression
-talent_tiny mfeat_fourier 1600 77 classification
-talent_tiny mv 32614 11 regression
-talent_tiny okcupid_stem 21341 14 classification
-talent_tiny online_shoppers 9864 15 classification
-talent_tiny optdigits 4496 65 classification
-talent_tiny page_blocks 4378 11 classification
-talent_tiny pc3 1250 38 classification
-talent_tiny pendigits 8793 17 classification
-talent_tiny pol 8065 27 classification
-talent_tiny pol_reg 12000 49 regression
-talent_tiny rl 3976 13 classification
-talent_tiny satimage 5144 37 classification
-talent_tiny segment 1848 18 classification
-talent_tiny socmob 924 6 regression
-talent_tiny splice 2552 61 classification
-talent_tiny sylvine 4099 21 classification
-talent_tiny thyroid_dis 2240 27 classification
-talent_tiny tic_tac_toe 31492 10 classification
-talent_tiny vulnonevul 4553 17 regression
-talent_tiny waterstress 950 23 regression
-talent_tiny waveform_5000 4000 41 classification
-talent_tiny website_phishing 1082 10 classification
-talent_tiny wine 2043 5 classification
-talent_tiny wine_quality_white 3918 12 classification
+We investigate the dataset-size dependent rank performance across all 203 datasets in Appendix LABEL:app:additional-results-data-size, further highlighting these challenges.
+Additional results, including additional baselines, CD diagrams, as well as 1 vs 1 win ratios and p-values, are provided in Appendix LABEL:app:additional-results-extended.
+Furthermore, we refer to Appendix LABEL:app:runtime-analysis for an analysis of the model runtime in comparison to other baselines.
+Table 1: Performance comparison across all evaluated benchmarks, depicting mean accuracy (Acc) for classification and (soft-clipped) $R^{2}$ score for regression tasks, in percent.
+Missing values, due to architectural limitations or failed evaluations, are denoted as N/A and excluded from the rank calculations.
+Models are sorted according to their ranking on the CARTE benchmark.